### v2

# Enhancing Financial Time-Series Forecasting with Retrieval-Augmented Large Language Models

Abstract.
Stock movement prediction, a critical task in financial time-series forecasting, relies on identifying and retrieving key influencing factors from vast and complex datasets. However, traditional text-trained or numeric similarity-based retrieval methods often struggle to handle the intricacies of financial data. To address this, we propose the first retrieval-augmented generation (RAG) framework specifically designed for financial time-series forecasting. Our framework incorporates three key innovations: a fine-tuned 1B large language model (StockLLM) as its backbone, a novel candidate selection method enhanced by LLM feedback, and a training objective that maximizes the similarity between queries and historically significant sequences. These advancements enable our retriever, FinSeer, to uncover meaningful patterns while effectively minimizing noise in complex financial datasets. To support robust evaluation, we also construct new datasets that integrate financial indicators and historical stock prices. Experimental results demonstrate that our RAG framework outperforms both the baseline StockLLM and random retrieval methods, showcasing its effectiveness. FinSeer, as the retriever, achieves an 8% higher accuracy on the BIGDATA22 benchmark and retrieves more impactful sequences compared to existing retrieval methods. This work highlights the importance of tailored retrieval models in financial forecasting and provides a novel, scalable framework for future research in the field.
Retrieval-augmented generation, Financial Time-series forecasting, Stock Movement Prediction, Large Language Model
copyright: noneccs: Information systems Top-k retrieval in databasesccs: Information systems Similarity measuresccs: Information systems Novelty in information retrieval
1. Introduction
Financial time-series forecasting plays a pivotal role in ensuring market stability and efficiency, directly impacting critical areas such as investment strategies, risk management, and the formulation of economic policies (Fama and French, 2000). One of the most prominent tasks in this domain is stock movement prediction (Xie et al., 2023a, 2024b, 2024a), which focuses on forecasting the direction of price changes, i.e. whether a stock will rise or fall. Accurate predictions of financial metrics, including stock movements, interest rates, and economic indicators, are essential for a wide range of stakeholders, from individual investors to large financial institutions. However, the inherent complexity and volatility of financial markets pose significant challenges to forecasting, necessitating the use of advanced methodologies to effectively analyze vast and often noisy datasets.
Traditional stock movement prediction methods (Feng et al., 2018; Qin et al., 2017) relied on analyzing sequences of stock prices and financial indicators [^1] Financial indicators (Jansen, 2020) are professional market analysis metrics derived from price data, where their values or crossover points signal fluctuations in the stock market, indicating rising or falling trends. However, relying on one or two indicators can introduce bias, and if an indicator only provides meaningful information at crossover points, it may fail to offer valuable insights during periods without crossings. to identify patterns, but they often failed to capture the semantic connections between these sequences, a task at which large language models (LLMs) excel (Huang and Chang, 2022). However, LLM-based methods (Xie et al., 2023b, 2024a, a) have primarily shifted focus to leveraging textual data such as news and tweets, often using only the past several days’ closing prices as reference while overlooking the rich insights embedded in historical time-series data (Bustos and Pomares-Quimbaya, 2020). This highlights the need to effectively integrate financial time-series data with LLMs. Directly working with such data presents a challenge due to its vast scale, encompassing both the breadth of influencing variables and the depth of historical trends. To address this, a retrieval-based approach is necessary to efficiently sift through extensive time-series data, extracting meaningful information to enhance LLMs’ ability to deliver more accurate and robust stock movement predictions.
Although retrieval-augmented generation (RAG) methods have been applied for various tasks (Joshi et al., 2024; He et al., 2024; Cui et al., 2024), it is challenging to directly apply them to financial time-series retrieval. First, most embedding-based retrieval methods (retrievers) (Su et al., 2023; Li and Li, 2023; Xiao et al., 2023; Zhang et al., 2023; Wang et al., 2023) are trained on textual data and struggle to process numeric time-series data, which lacks explicit semantic information. Second, distance-based retrieval methods like Dynamic Time Warping (DTW) (Yang et al., 2024) focus on numeric similarities and fail to capture deeper semantic relationships or contextual nuances essential for stock movement prediction. These limitations underscore the need for a more tailored retrieval framework designed for financial time-series data.
[…]
The framework of FinSeer, illustrating its components and workflow.
Figure 1. Overview of our time-series RAG framework.
To address these challenges, we introduce a retrieval-augmented framework for stock movement prediction, featuring the first dedicated retriever for financial time-series forecasting, Financial Time-Series Retriever (FinSeer). Unlike previous RAG methods, which rely on pre-trained encoders or distance-based metrics, our framework is the first to train a dedicated retriever tailored for continuous and complex temporal sequences. The whole retrieval framework is shown in Figure 1.
To address the limitations of existing datasets, which often contain only price data, and to better align with professional financial analysis, we construct new datasets by integrating 20 groups of financial indicators. These indicators provide critical insights into market behavior that stock prices alone cannot capture [^2] For instance, the overbought area in the KDJ indicator group (Wu and Diao, 2015), where $K>80$, $D>70$, and $J>90$, signals a potential falling trend.. We begin by selecting high-trade-volume U.S. stocks, and the indicators are chosen through domain expert consultation and mutual information analysis. These indicators are then serialized and divided into train, valid and test sets. By training the retriever on datasets enriched with diverse financial indicators, our retrieval framework incorporates financial domain expertise, enabling it to interpret the implications of trending signals that uncover deeper patterns beyond surface-level trend similarities. The testing set, designed to mimic real-world economic analysis, ensures a robust evaluation of model performance in practical financial scenarios.
To train our retriever to identify semantically related sequences and align the retriever’s priorities with those of the LLM, we introduce a novel retrieval mechanism with three key innovations: a new backbone LLM, a novel method for selecting positive and negative candidates, and a new training objective. First, to activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA3.2-1B Instruct) (Grattafiori et al., 2024), creating StockLLM as the backbone model of our retriever. By using a smaller LLM, we establish a more challenging experimental setup, ensuring performance improvements are attributable to FinSeer rather than the LLM’s capacity. Next, StockLLM identifies time-series segments that enhance the generative process, feedbacks from StockLLM are used to select the most beneficial sequences that could lead the model to make the correct decision as positive candidates and the least beneficial as negative candidates. This could bridge the gap between the retriever and the LLM and ensure that retrieved data aligns with the LLM’s forecasting priorities. Finally, the training objective, inspired by Zhang et al. (2023), ensures the retriever distinguishing historically significant sequences (positive candidates) from noisy sequences (negative candidates) by maximizing the similarity between the query and candidate sequences. This prioritizes meaningful patterns while minimizing irrelevant data, outperforming traditional methods that struggle to separate signal from noise in complex financial time series.
Based on FinSeer, the inference process of our RAG framework consists of three steps: (1) using FinSeer to retrieve relevant sequences from the candidate pool, (2) incorporating these sequences into StockLLM’s input context, and (3) combining this enriched context with the original query to perform forecasting. This approach effectively integrates salient historical patterns with current data, improving prediction accuracy.
To thoroughly evaluate the performance of our RAG framework and the retriever, we rely solely on temporal information, excluding additional textual data like news. For the RAG framework evaluation, StockLLM serves as the backbone model, integrated with different retrieval methods for comparison. Experimental results demonstrate that using our RAG framework outperforms bare StockLLM, indicating that the RAG framework enhances LLM performance by incorporating relevant time-series data as financial context. However, RAG with retrieval models trained on text data does not always improve performance compared to bare StockLLM. In some cases, performance even declines. This suggests that these retrieval models cannot effectively identify beneficial time-series data that could enhance the model’s predictions, which highlights a fundamental gap between text-based retrieval and time-series data. Furthermore, RAG framework with FinSeer consistently surpasses all other retrieval models, such as Instructor (Su et al., 2023), BGE (BAAI General Embedding) (Xiao et al., 2023), LLM-Embedder (Zhang et al., 2023), UAE (Li and Li, 2023), and E5-mistral-7b-instruct (Wang et al., 2023), demonstrating the effectiveness of our trained FinSeer in identifying relevant time-series financial data to enhance stock movement prediction. Further analysis of retrieval results reveals that FinSeer uniquely identifies relevant financial indicators that enhance LLM performance, while other retrievers predominantly select sequences based on superficial similarities, such as close price and adjusted close price. This behavior underscores a critical limitation of text-trained models: they focus on surface-level patterns and fail to discern truly valuable time-series information. In contrast, FinSeer effectively bridges this gap, demonstrating its ability to uncover underlying relevance in financial data.
Our contributions are summarized as follows:
  1. (1)
    We propose the first RAG framework for financial time-series forecasting, in which the retrieval mechanism features three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a novel candidate selection method leveraging LLM feedback to identify beneficial sequences, and a training objective that maximizes similarity between queries and historically significant sequences, enabling the retriever to uncover meaningful patterns while minimizing noise in complex financial data.
  2. (2)
    We construct new training and testing datasets to address the limitations of existing ones containing only price data, and better align with real-world economic analysis. Besides price data, these datasets integrate financial indicators that signal upward or downward trends, carefully selected through domain expert consultation and mutual information analysis. By using new training datasets, we train the retriever to uncover deeper patterns beyond surface-level trends during training. While the testing set, designed to mimic real-world scenarios, ensures robust evaluation of model performance in practical financial applications.
  3. (3)
    We propose a new retriever, FinSeer, specifically designed for time-series forecasting, trained on newly constructed datasets that integrate financial indicators alongside historical stock prices. By leveraging LLM feedback and prioritizing historically significant sequences, FinSeer uncovers deeper patterns beyond surface-level trends, setting it apart from traditional methods.
  4. (4)
    Experimental results demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, highlighting its effectiveness, while our retriever, FinSeer, surpasses existing RAG methods in stock movement prediction across three datasets, achieving an 8% higher accuracy on BIGDATA22 compared to a general-purpose LLM-feedback-based retriever. Furthermore, FinSeer retrieves more dynamic and impactful sequences, showcasing its ability to identify and leverage the most relevant patterns for forecasting.
2. Problem Definition and Goals
Retrieval-augmented financial time-series forecasting (Sezer et al., 2020) involves predicting future values or trends ($G$) based on a given query sequence ($q$) and a set of retrieved historical sequences ($c$). These sequences are collected over time at regular intervals. In the retrieval process, the goal of the retrieval model ($R$) is to efficiently identify and extract the most useful historical sequences from a large pool of candidates. By providing relevant context, the retrieval model enhances the forecasting model’s ability to make accurate and reliable predictions.
Stock Movement Prediction Task. In the specific task of stock movement prediction (Xu and Cohen, 2018), the problem is framed as a binary classification task: predicting whether a stock’s price will rise or fall on the next trading day. Given a query sequence $q$, which represents the stock’s price over the previous $t$ days, the model retrieves relevant sequences as context and predicts the stock’s movement $M_{q,d}$ for the next trading day $d$. Table 1 defines the major symbols used in this paper.
Table 1. The definition of symbols.
General Time-series Symbol Stock Movement Prediction Symbol Definitions
$q$ $q=\{q_{d-t},...,q_{d-1}\}$ The query time-series data. In stock movement prediction, $q$ refers to the query stock price sequence of length $t$, containing stock price data from trading day $d-t$ to trading day $d-1$.
$G(q)$ $G(q,d)\in\{\textit{rise},\textit{fall}\}$ The final output $G$ given the query $q$. In stock movement prediction, $G(q,d)$ shows the generation $G$ of the query stock $q$ on the query trading day $d$, belonging to rise or fall.
$P(c)=LLM(O|q,c)$ $P(c)=LLM(M_{d}|q,c)$ The probability $P$ of the LLM to generate an accurate output $O$ given the query sequence $q$ and a candidate sequence $c$. In stock movement prediction, $P(c)$ refers to generating the accurate movement $M$ on the query trading day $d$.
=$\mathbb{C_{P}}=\{c_{i}\mid i=1,\ldots,k\}$ The set of top-k retrieved sequences as positive examples, where $P(c_{i})\geq P(c_{i+1})$.
=$\mathbb{C_{N}}=\{c_{i}\mid i=k+1,\ldots,n\}$ The set of negative retrieved sequences, where $P(c_{i})\geq P(c_{i+1})$.
$w_{i}$ The soft weight of the $i^{th}$ retrieved sequences, where $w_{i}=P(c_{i}),i=1,...,k$.
$R$ The retrieve model.
Rise/Fall Threshold Settings. To classify daily movements as rise or fall, we first calculate returns $R_{t}$, which represents the percentage change in the closing price over consecutive days.
[…]
Following Yoo et al. (2021) and  Soun et al. (2022), we classify the movement as rise if return exceeds 0.55, fall if it is below -0.5. In line with previous experimental settings (Xie et al., 2023b, 2024a), we do not evaluate freeze cases. However, we include sequences with $R_{t}\in[-0.5,0.55]$ as freeze candidates in the candidate pool, ensuring a diverse and comprehensive set of historical data for context.
$M_{q,d}&=\begin{cases}\text{rise},&R_{t}>0.55\\ \text{fall},&R_{t}<-0.5\\ \text{freeze},&-0.5\leq R_{t}\leq 0.55\end{cases}\end{aligned}$
Rationale for Unbalanced Thresholds. The asymmetrical thresholds for classifying daily movements (0.55 for rises and -0.5 for falls) reflect the inherent dynamics of stock market behavior. Stock prices typically rise gradually due to sustained investor optimism but fall sharply during panic selling or profit-taking. The stricter rise threshold prevents minor upward fluctuations from being misclassified as significant increases, while the more lenient fall threshold ensures meaningful downward trends are captured. This approach aligns with market realities, improving the reliability of movement classifications.
3. The RAG Framework
To identify valuable sequences from vast amounts of stock data and conduct stock movement prediction, we propose a novel RAG framework with new datasets and a novel retrieval mechanism with a new backbone LLM. First, we construct three datasets that integrate financial indicators alongside historical stock prices, addressing the limitations of existing datasets and better aligning with real-world economic analysis. Second, we introduce StockLLM, a fine-tuned 1B parameter LLM, as the backbone to ensure instruction-following capabilities and establish a challenging experimental setup. Finally, we design a novel retrieval mechanism that leverages LLM feedback to select positive and negative candidates and employs a new training objective to prioritize historically significant sequences. This comprehensive framework bridges the gap between retrieval models and LLMs, enhancing the time-series forecasting process.
3.1. Dataset Construction
To address the limitations of existing datasets and better align with real-world economic analysis, we construct new datasets by integrating financial indicators alongside historical stock prices, enabling deeper insights into market behavior that stock prices alone cannot capture. This section details stock selection, raw data collection (Collection of Basic Price Indicators) and financial indicator selection process(Calculation and Selection of Financial Indicators). Specifically, our datasets are built using high-trade-volume U.S. stocks, while the indicators are selected through domain expert consultation and mutual information analysis.
Stock Selection. To construct our datasets, we select high-trade-volume U.S. stocks across three periods: 2014-2015, 2017-2018, and 2022-2023. The first two periods align with two benchmark datasets: for 2014-2015, we use the same stocks as the ACL18 dataset (Xu and Cohen, 2018), and for 2017-2018, we use the same stocks as the BIGDATA22 dataset (Soun et al., 2022). To incorporate more recent data, we manually select high-trade-volume stocks from 2022-2023 to create the STOCK23 dataset. This ensures our training and evaluation reflect recent market conditions and provide a robust benchmark for stock movement prediction.
Collection of Basic Price Indicators. Stock data is collected using the Yahoo Finance API (Xu and Berkely, 2014), which provides basic price indicators such as the opening price, highest price, lowest price, adjusted closing price, and trading volume for each trading day. The daily price movement is then calculated as the reference answer for model training and evaluation.
We then divide the stock data into training, validation, and test sets. To ensure temporal integrity and prevent information leakage, we partition the dataset by stocks rather than by time. This approach maintains consistent time ranges across splits while ensuring that stocks in the validation and test sets are entirely unseen during training. Unlike time-based splits, stock-based partitioning provides a more robust evaluation of the model’s ability to generalize to new stocks, better reflecting real-world scenarios where models are applied to unseen entities. Table 2 summarizes the stock counts for each split in our datasets.
Table 2. Dataset statistics.
Dataset Stock amount Trading dates
train valid test all sequences query sequences
ACL18 (Xu and Cohen, 2018) 33 5 33 2014.06.02-2015.12.31 2015.06.03-2015.12.31
BIGDATA22 (Soun et al., 2022) 22 3 22 2019.04.01-2020.12.31 2020.04.09-2020.12.31
STOCK23 24 3 24 2022.01.03-2023.12.31 2023.01.03-2023.12.31
Calculation and Selection of Financial Indicators. In our datasets, we integrate financial indicators to provide critical insights into market behavior that stock prices alone cannot capture. Following analyzing methods in the book Machine learning for trading (Jansen, 2020), we conduct an analysis using query sequences in all training sets to explore the relationship between commonly used financial indicators and returns. To measure the relationship between these indicators and returns, we utilize Mutual Information (MI), a non-linear measure of dependency that quantifies the amount of information one variable contains about another. Specifically, the MI between an indicator $X$ and returns $R_{t}$ is calculated as:
[…]
where $p(x,r)$ represents joint probability density function of $X$ (indicator) and $R_{t}$ (returns), $p(x)$ shows marginal probability density function of $X$, and $p(r)$ is marginal probability density function of $R_{t}$.
We compute MI scores for each indicator and select the top-20 indicator groups with the highest scores as our candidates, with the detailed score ranking provided in Appendix B.1. Additionally, an example of a data piece from our datastore is illustrated in Appendix A.1.
3.2. Sequence Serialization
Since stock movement prediction depends on the changes in related features rather than their exact values, we serialize stock prices and financial indicators into a time-series format. We use JSON to represent these sequences, as it has been demonstrated to effectively support LLMs in interpreting time-series data (Fang et al., 2024; Singha et al., 2023; Yin et al., 2023).
Query Sequence Serialization We select query sequences where the query date is at least one year after the start date of the corresponding split in the dataset. This ensures that each query has enough candidate sequences for retrieval. For example, in the ACL18 dataset, which spans from 2014-06-02 to 2015-12-31, the query sequences are defined as those from 2015-06-03 to 2015-12-31.
Each query sequence contains the adjusted close price from the previous five trading days along with basic stock information. The indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days. For example, when inquerying about stock MO on 2015-06-02, the query sequences contains stock name, query date, last five trading dates and their corresponding adjusted close prices. The serialized sequence is shown below:
{ "query_stock": "MO",
"query_date": "2015-06-02",
"recent_date_list": ["2015-05-26", "2015-05-27", "2015-05-28", "2015-05-29", "2015-06-01"],
"adjusted_close_list": [29.669, 29.9872, 29.8657, 29.6227, 29.6227]}
Candidate Sequence Serialization As new information becomes available, the candidate pool dynamically incorporates sequences from the most recent trading days. For instance, when the query date is 2015-06-03, the candidate pool includes sequences from 2014-06-02 to 2015-06-02. When the query date advances to 2015-06-04, the sequence for 2015-06-03 is added to the candidate pool without the need for additional training. This demonstrates the scalability of our method.
For each candidate stock on a specific date, we select 6 basic price indicators and 20 groups of financial indicators. Each candidate sequence includes a typical financial indicator along with basic stock information. As the same as query serialization, the indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days. For example, a candidate represent stock MO on date 2014-07-02, with highest price as its indicator. The sequence includes basic information containing stock name, date, movement, last five trading dates and their corresponding highest prices. The serialized sequence is shown below:
{ "candidate_stock": "MO",
"candidate_date": "2014-07-02",
"candidate_movement": "freeze",
"recent_date_list": ["2014-06-25", "2014-06-26", "2014-06-27", "2014-06-30", "2014-07-01"],
"high_list": [42.2, 42.0, 41.86, 42.28, 42.0]}
3.3. Retriever Training
To identify valuable information from vast amounts of sequences and uncover semantic relationships between them, we introduce a novel retrieval mechanism featuring three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a candidate selection method leveraging LLM feedback to identify beneficial sequences and a training objective that maximizes similarity between queries and historically significant sequences, enabling the trained retriever to uncover meaningful patterns while minimizing noise in complex financial data. While inspired by (Zhang et al., 2023), our approach is uniquely tailored to financial time-series data, addressing the challenges of relevance assessment and alignment in a domain where traditional retrieval methods often fall short. This section details the training process of our framework, including the training of the LLM backbone, candidate scoring and selection, and the training of the retriever, which contains the training objective and knowledge distillation.
The LLM Backbone. To activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA 3.2-1B-Instruct) (Grattafiori et al., 2024) using the LoRA technique for efficient low-rank adaptation, resulting in StockLLM. By intentionally using a smaller backbone model, we establish a more challenging experimental setup, ensuring that performance improvements are attributable to FinSeer’s retrieval capabilities rather than the LLM’s capacity. The fine-tuning process is implemented using the LlamaFactory framework (Zheng et al., 2024), with the following configuration: a learning rate of 5e-5, a cosine scheduler, gradient accumulation over 8 steps, mixed-precision (fp16) training, and 5 epochs of training with regular evaluation to log metrics and select the best-performing checkpoint. This setup demonstrates the robustness of our framework in extracting meaningful insights even under constrained model size, highlighting the effectiveness of FinSeer in enhancing financial forecasting tasks.
Candidate Scoring. To determine whether a candidate sequence assists in predicting the movement of the query, we use LLM feedback to score each candidate. Specifically, for a given query $q$, we integrate the query sequence and each candidate sequence $c_{i}$ from the candidate pool as concurrent inputs to the LLM. The LLM outputs logits, which are unnormalized scores representing the model’s confidence for each possible class (e.g., ”rise” or ”fall”). These logits are transformed into probabilities $P(c)$ using the softmax function:
(4) $P(c)=\frac{e^{z_{c}}}{\sum_{j}e^{z_{j}}},$
where $z_{c}$ is the logit for the correct class (e.g., ”rise” if the true movement is upward) and $z_{j}$ represents the logits for all possible classes. The resulting probability $P(c)$ serves as the score for the candidate $c_{i}$ with respect to the query $q$.
Candidate Selection. We rank the candidate sequences in descending order based on their scores $P(c)$. The top-1 sequence is selected as a positive candidate, while the bottom 15 sequences are chosen as negative candidates. The sets of selected positive and negative sequences are denoted as $\mathbb{C_{P}}$ and $\mathbb{C_{N}}$, respectively.
Training Objective. Our retriever $R(q)$ is designed to intelligently distinguish between historically significant sequences $\mathbb{C_{P}}$ and noisy sequences $\mathbb{C_{N}}$. The training objective is to ensure that $R(q)$ prioritizes sequences from $\mathbb{C_{P}}$ while minimizing attention to those from $\mathbb{C_{N}}$. This is achieved by maximizing a similarity measure $sup(q,s)$ between the query sequence $q$ and candidate sequences $s$. Mathematically, the retriever’s objective is formulated as:
(5) $R(q)={\arg\max}_{s\in\mathbb{C_{P}}\cup\mathbb{C_{N}}}sup(q,s).$
By focusing on sequences that maximize $sup(q,s)$, the retriever ensures that the most informative and contextually relevant historical sequences are identified.
Knowledge Distillation. To leverage the scoring derived from the LLM, we employ knowledge distillation, which transfers knowledge from the teacher model (LLM) to the student model (retriever) by mimicking the teacher’s output distribution. This approach effectively captures nuanced patterns and predictions from the LLM. Specifically, we minimize the Kullback-Leibler (KL) divergence between the candidate distributions computed using the LLM’s rewards and those predicted by the embedding model. For each query $q$ and its candidate list {$\{\mathbb{C_{P}},\mathbb{C_{N}}\}$, we derive the LLM’s rewards for the candidates, denoted as $\{P(c_{i}),i=1,...,n\}$. To make these rewards suitable for distillation, we normalize them using a softmax function with temperature $\alpha$:
(6) $w_{i}=\text{softmax}_{R}\left(\frac{P(c_{i})}{\alpha}\right).$
The KL divergence is then computed as follows:
(7) $\min\sum_{c}-w_{i}\times\log\left(\frac{\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{c_{i}}\right\rangle/\tau\right)}{\sum_{{c}^{\prime}\in\mathbb{C}}\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{{c}^{\prime}}\right\rangle/\tau\right)}\right),$
where $\boldsymbol{e}_{q}$ and $\boldsymbol{e}_{c_{i}}$ are the embeddings of the query $q$ and candidate $c_{i}$, respectively, and $\tau$ is a temperature parameter. This loss function optimizes the similarity between the query embedding and the embeddings of the top-ranked candidates, enhancing the retriever’s ability to accurately predict stock price movements.
3.4. Inference
The inference process of our RAG framework, built on FinSeer, consists of three key steps to enhance stock movement prediction. First, FinSeer retrieves relevant sequences from the candidate pool, leveraging its ability to identify historically significant patterns. Second, these retrieved sequences are incorporated into StockLLM’s input context, enriching the model’s understanding of temporal dynamics. Finally, the enriched context is combined with the original query to generate forecasts. By integrating salient historical patterns with current data, this approach bridges the gap between past trends and present conditions, significantly improving prediction accuracy and robustness in real-world financial scenarios.
4. Experiment
4.1. Experimental Settings
Datasets. We evaluate the effectiveness of our RAG framework on the test sets of the three datasets described in Table 2, with ACL18 containing 2,876 query sequences, BIGDATA22 containing 2,868 queries, and STOCK23 containing 4,128 queries. These thousand-scale queries help mitigate random bias, ensuring a robust and reliable evaluation of model performance. To ensure a comprehensive evaluation, for each query sequence, we include all sequences containing financial indicators across all stocks in the test set (not limited to the same stock), with data available up to the query date, as potential candidates. No additional restrictions are imposed, enabling a robust assessment of the models’ performance in real-world financial forecasting scenarios.
Baselines. To evaluate the efficiency of our RAG framework, we choose bare StockLLM and random retrieval as our baselines. To evaluate our retriever FinSeer, we select five competitive RAG models from the top of the MTEB English Leaderboard as baselines: (1) Instructor-large (Su et al., 2023): A 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors. (2) UAE-large-v1 (Li and Li, 2023): A 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors. (3) BGE-large-en-v1.5 (Xiao et al., 2023): A 335M general embedder pre-trained with RetroMAE (Xiao et al., 2022), encoding sequences into 1024-dimensional tensors. (4) LLM-Embedder (Zhang et al., 2023): A 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs. It encodes sequences into 768-dimensional tensors. (5) E5-mistral-7b-instruct (Wang et al., 2023): A 7111M embedder initialized from Mistral-7B-v0.1 (Jiang et al., 2023) and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
Evaluation Metrics. We employ Accuracy (ACC) and Matthews Correlation Coefficient (MCC) (Matthews, 1975) to assess the performance of FinSeer and the baseline models on the stock movement prediction task. These metrics evaluate the performance of stock movement prediction based on the distribution of positive and negative samples.
4.2. Main Results
Table 3 summarizes the performance of stock movement prediction across the three datasets.
Table 3. Results of stock movement predictions using LLMs and retrieval models. The asterisk (*) indicates the LLM employed while using retrieval models.
Retrieving Methods ACL18 BIGDATA22 STOCK23
(+ StockLLM) ACC MCC ACC MCC ACC MCC
w/o Retrieval 0.496 -0.007 0.497 -0.008 0.509 0.021
Random Retrieve 0.481 -0.036 0.521 0.045 0.496 -0.004
DTW 0.518 0.045 0.505 0.021 0.492 -0.007
Instructor 0.496 -0.008 0.497 -0.006 0.505 0.010
UAE 0.482 -0.036 0.495 -0.008 0.494 -0.009
E5 0.510 0.028 0.503 0.008 0.499 -0.001
BGE 0.489 -0.019 0.487 -0.026 0.488 -0.014
LLM Embedder 0.499 0.000 0.459 -0.083 0.503 0.007
FinSeer 0.521 0.042 0.538 0.079 0.542 0.085
We analyze our framework and retriever separately to evaluate their contributions. First, our RAG framework outperforms both bare StockLLM and random retrieval across all datasets, demonstrating its effectiveness in enhancing stock movement prediction. Second, among all evaluated retrieval methods, FinSeer consistently achieves the best performance when integrated with StockLLM, showcasing its ability to retrieve valuable sequences that improve forecasting accuracy. Notably, FinSeer is the only retriever that achieves stable improvements over bare StockLLM on all datasets, while other methods exhibit inconsistent results, sometimes enhancing predictions and other times degrading them. Furthermore, our experiments reveal that incorporating a retrieval process does not always guarantee performance improvements, as retrieval errors can mislead the LLM and harm predictions. This underscores the critical need for retrieval models that align closely with the LLM’s objectives, a challenge that FinSeer effectively addresses through its tailored design.
Figure 2. Indicator occurrences of different RAG models on ACL18 dataset.
The minimal performance difference between the instruction-finetuned retriever (Instructor) and the no-retrieval baselines highlights the inherent challenge of time-series retrieval. Unlike text retrieval, time-series retrieval does not solely rely on task understanding, as candidate sequences often appear similar, making it difficult to distinguish valuable ones. The angle-optimized retriever (UAE) shows inconsistent performance across datasets, as its ability to differentiate similarity between query and candidate sequences does not fully align with the LLM’s perception of importance. Similarly, LLM Embedder, our backbone model trained with LLM feedback, struggles to generalize to time-series retrieval tasks, further underscoring the complexity of the problem.
The best-performing baseline is E5, which achieves competitive results despite its significantly larger model size and embedding dimensions (five times larger than FinSeer). However, FinSeer outperforms E5 on all datasets. This demonstrates that FinSeer successfully learns the preferences of LLMs and effectively enhances their time-series forecasting ability through retrieval, even under challenging conditions.
In summary, while RAG methods do not guarantee performance improvements due to the risk of retrieval errors, FinSeer stands out by consistently retrieving high-quality sequences that align with the LLM’s forecasting objectives, leading to superior results compared to other retrievers.
4.3. Ablation Study
Indicator Occurrences In this part, we analyze what indicators all RAG models retrieve. We calculate indicator occurrences on ACL18 dataset and the results are shown in Figure 2.
As shown in the figure, FinSeer is the only model that successfully extracts a diverse and comprehensive set of indicators while achieving superior performance. This clearly demonstrates its advanced temporal retrieval capabilities. Specifically, while other models like LLM Embedder and Instructor predominantly focus on basic indicators such as close price and adjusted close price, FinSeer effectively identifies and retrieves a wide range of technical indicators, including kdj crossover, MACD Histogram, Bollinger Bands, and various alpha factors. This richer set of retrieved indicators provides FinSeer with more comprehensive auxiliary information, enabling more accurate and reliable predictions.
Candidate Movement Correlation In this section, we investigate whether the LLM relies on the movement trends of retrieved sequences to make predictions. To this end, we compute the correlation between the movements of retrieved sequences and the LLM’s generated results. The detailed calculation process is provided in Appendix B.2.
The results in Table 4 reveal no significant correlation between the movement direction (rise or fall) of the retrieved candidate sequences and the final predictions of StockLLM. This finding indicates that StockLLM does not simply mirror the movement trends of the retrieved sequences but instead analyzes their specific content to infer the query’s movement direction. This ability highlights the LLM’s capacity to extract meaningful insights from complex time-series data.
Table 4. Correlation between the movements of retrieved sequences and LLM’s generated results.
Dataset RAG Model Correlation P-Value
ACL18 Instructor -0.098 0.000
UAE -0.068 0.000
E5 -0.108 0.000
BGE 0.175 0.000
LLM Embedder 0.014 0.457
FinSeer 0.054 0.004
BigData22 Instructor -0.072 0.000
UAE 0.009 0.633
E5 -0.040 0.032
BGE 0.093 0.000
LLM Embedder 0.060 0.001
FinSeer 0.077 0.000
STOCK23 Instructor -0.114 0.000
UAE -0.093 0.000
E5 -0.064 0.000
BGE 0.131 0.000
LLM Embedder 0.015 0.350
FinSeer 0.032 0.041
This observation underscores the critical role of retrieval quality in RAG models, as the relevance and informativeness of retrieved sequences directly influence prediction performance. Notably, StockLLM is a relatively small 1B parameter LLM, making this a particularly challenging setting. Despite this, our time-series-tailored RAG model, FinSeer, successfully enhances StockLLM’s performance by retrieving rich and relevant time-sensitive information. This demonstrates the effectiveness of our approach in improving prediction accuracy and showcases its potential for financial time-series forecasting.
4.4. Case Study
This case study illustrates the critical importance of alignment between the retriever and the LLM’s forecasting preferences in financial time-series analysis. We examine the stock XOM on 2015-06-25 from the ACL18 dataset, where the adjusted close price exhibited a pronounced downward trend. The query sequence is as follows:
{ "query_stock": "XOM",
"query_date": "2015-06-25",
"recent_date_list": ["2015-06-18", "2015-06-19", "2015-06-22", "2015-06-23", "2015-06-24"],
"adjusted_close_list": [58.0813, 57.8979, 57.8707, 57.8027, 57.5377]}
While multiple retrievers were evaluated, only FinSeer successfully enabled StockLLM to predict the correct movement as a fall. Specifically, FinSeer retrieved five diverse indicators—close price, adjusted close price, alpha021, alpha054, and the highest price—providing a comprehensive view of the stock’s behavior. Alpha021 identifies trends based on short- and long-term price averages and volume conditions, while alpha054 combines price and volume rankings to assess performance within a specific time window. These indicators allowed StockLLM to accurately assess whether the downward trend would persist or reverse, demonstrating the value of retrieving contextually relevant and diverse features.
In contrast, other retrievers, such as Instructor, BGE, LLM Embedder, and E5, extracted sequences dominated by close or adjusted close prices, all reflecting similar downward trends. While these sequences aligned with the current trend, they failed to provide actionable insights for forecasting future movements, leading StockLLM to misinterpret them as noise and incorrectly predict a rise. Similarly, UAE retrieved sequences indicating overbought and oversold conditions, including three rise and two freeze trends. Although overbought signals often suggest a potential downturn, the retrieved sequences themselves exhibited rising or frozen trends, confusing StockLLM and resulting in an erroneous prediction. This case study underscores the superiority of FinSeer in retrieving meaningful and diverse indicators that align with the LLM’s forecasting logic, enabling more accurate and reliable predictions.
4.5. Indicator Sequence Visualization
To intuitively examine how FinSeer embeds indicator sequences, we use LargeVis (Tang et al., 2016) to reduce the embeddings to two dimensions. In Figure 3, we visualize [^3] https://medviz.org/app/ the vector space of nine indicators, including adjusted close price, open price, high price, low price, close price, volume, movement, returns, and WAP. The results show that the green points, representing movement information, are well-clustered, indicating that FinSeer effectively captures meaningful patterns. However, this also suggests that deeper relationships between stock movements and indicators remain to be explored, highlighting the potential for further research in this area.
Figure 3. Visualization of indicator sequence embeddings on ACL18 dataset.
[…]
5. Related Work
5.1. Stock Movement Prediction
Non-LLM Methods. Traditional approaches to stock movement prediction have focused on various aspects of financial data. One prominent category of methods analyzes stock price sequences and their corresponding technical indicators (Qin et al., 2017; Feng et al., 2018) to identify patterns in historical data for predicting future movements. However, due to the complexity of factors influencing stock prices, subsequent methods have incorporated additional contextual information, such as news articles (Ding et al., 2015; Liu et al., 2018) or social media posts (Xu and Cohen, 2018; Wu et al., 2018). Despite efforts, these methods are highly susceptible to noise and struggle to analyze the vast and diverse nature of financial information effectively.
LLM-based Methods. Recent studies have explored using LLMs for financial prediction tasks, either by fine-tuning open-source models or prompting advanced models like GPT-4. However, even state-of-the-art models, including GPT-4, have achieved only random-guessing-level accuracy in stock movement prediction (Xie et al., 2023b, 2024a, a). This highlights the inherent challenges in identifying and analyzing meaningful patterns in a domain as volatile and multifaceted as stock market prediction.
5.2. Time-series forecasting with LLMs
Non-retrieval Methods. To enhance the performance of LLMs in time-series forecasting, existing methods have primarily focused on aligning temporal and textual data, either by transforming time-series into textual formats or by encoding both modalities into a unified vector space. For instance, a study (Jin et al., 2023) reprograms time-series data into textual representations suitable for LLMs, enhancing prediction accuracy via declarative prompts. Similarly, some other studies (Yu et al., 2023; Liu et al., 2024) explore cross-modal alignment: the former applies LLMs to financial forecasting by integrating stock prices with news data, while the latter introduces a cross-modality framework to align time-series with text for improved predictive performance. While these advancements improve alignment, existing methods often struggle to process large-scale time-series data comprehensively due to LLM input limitations. As a result, effectively leveraging extensive historical data remains a challenge. This highlights the need for retrieval-augmented methods, a gap our approach directly addresses.
RAG Methods. With the rapid advancement of RAG techniques in various applications of LLMs, recent research has begun exploring RAG for time-series forecasting. For instance, TimeRAG (Yang et al., 2024) integrates RAG into time-series forecasting by combining Dynamic Time Warping (DTW) with LLMs to improve prediction accuracy. However, relying solely on numeric similarity is insufficient for financial time-series forecasting, as it fails to capture deeper semantic relationships. This underscores the need for a more targeted retrieval framework tailored to the complexities of financial data, a challenge our framework effectively addresses.
6. Conclusion
In this work, we introduce the first RAG framework for financial time-series forecasting, featuring a novel retrieval mechanism with three key innovations: a fine-tuned 1B parameter LLM (StockLLM), a candidate selection method leveraging LLM feedback, and a training objective that maximizes similarity between queries and historically significant sequences. We also construct new datasets integrating financial indicators and historical stock prices, selected through expert consultation and mutual information analysis, to train the retriever and ensure robust evaluation. Our retriever, FinSeer, leverages these datasets and LLM feedback to uncover deeper patterns beyond numeric similarities. Experiments demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, while FinSeer surpasses existing retrieval methods, achieving an 8% higher accuracy on BIGDATA22 and retrieving more impactful sequences. This work highlights the importance of tailored retrieval models in financial forecasting and provides a robust framework for future research.
Appendix A Data examples
A.1. Examples in the Datastore
An example in our datastore is shown below.
"stock_name":"ABBV",
"query_date":"2014-06-05",
"query_movement":"rise",
"open":54.549999,
"high":55.32,
"low":54.360001,
"close":55.299999,
"adj_close":36.980961,
"volume":4847300,
"MACD_Histogram":0.0621141602,
"macd_crossover":null,
"bollinger_bands":null,
"exceeding_upper":null,
"exceeding_lower":null,
"overbought_and_oversold_conditions":null,
"kdj_crossover":null,
"Returns":0.0131913094,
"VWAP":54.9933333333,
"alpha_smr":0.000549858,
"alpha_mom":0.0121661224,
"alpha002":null,
"alpha006":null,
"alpha009":-0.481476,
"alpha012":-0.481476,
"alpha021":1,
"alpha023":0.0,
"alpha024":-0.769035,
"alpha028":null,
"alpha032":null,
"alpha041":-0.1554335256,
"alpha046":-0.481476,
"alpha049":-0.481476,
"alpha051":-0.481476,
"alpha053":null,
"alpha054":126.4264005641,
"alpha101":-18.282056485
Appendix B Addtional Experimental Results
B.1. Score Calculation of Financial Indicators
We compute the MI scores for each indicator using nonparametric estimation methods provided by mutual_info_regression from the Scikit-learn library. Then we normalize them into range $(0,1)$. These scores provide insight into the strength of the dependency between each financial indicator and the forward return. Higher MI scores indicate stronger relationships, highlighting which indicators are most predictive of future price movements. We select the top-20 indicator groups with the highest MI scores in our candidates, as is shown in Figure 4.
Figure 4. Mutual Information(MI) scores between indicators and forward return.
B.2. Candidate Movement Correlation
For each query with retrieved candidates, we convert the movements of retrieved sequences and LLM’s prediction to numerical representations.
${M_{c_{j}},M_{q}}&=\begin{cases}\text{1},&\text{rise}\\ \text{0},&\text{freeze}\\ \text{-1},&\text{fall}\end{cases}\end{aligned}$
Then, we calculate the average value of the $M_{c_{i}}$ from the five retrieved sequences to represent the RAG-provided result, and the final prediction of the query is recorded as another numerical value:
[…]
Lastly, we adopt the Pearson correlation coefficient to compute the correlation:
[…]

### v3

@@
-Enhancing Financial Time-Series Forecasting with Retrieval-Augmented Large Language Models
-Abstract.
-Stock movement prediction, a critical task in financial time-series forecasting, relies on identifying and retrieving key influencing factors from vast and complex datasets.
-However, traditional text-trained or numeric similarity-based retrieval methods often struggle to handle the intricacies of financial data.
-To address this, we propose the first retrieval-augmented generation (RAG) framework specifically designed for financial time-series forecasting.
-Our framework incorporates three key innovations: a fine-tuned 1B large language model (StockLLM) as its backbone, a novel candidate selection method enhanced by LLM feedback, and a training objective that maximizes the similarity between queries and historically significant sequences.
-These advancements enable our retriever, FinSeer, to uncover meaningful patterns while effectively minimizing noise in complex financial datasets.
-To support robust evaluation, we also construct new datasets that integrate financial indicators and historical stock prices.
-Experimental results demonstrate that our RAG framework outperforms both the baseline StockLLM and random retrieval methods, showcasing its effectiveness.
-FinSeer, as the retriever, achieves an 8% higher accuracy on the BIGDATA22 benchmark and retrieves more impactful sequences compared to existing retrieval methods.
-This work highlights the importance of tailored retrieval models in financial forecasting and provides a novel, scalable framework for future research in the field.
-Retrieval-augmented generation, Financial Time-series forecasting, Stock Movement Prediction, Large Language Model
-copyright: noneccs: Information systems Top-k retrieval in databasesccs: Information systems Similarity measuresccs: Information systems Novelty in information retrieval
-1. Introduction
-Financial time-series forecasting plays a pivotal role in ensuring market stability and efficiency, directly impacting critical areas such as investment strategies, risk management, and the formulation of economic policies (Fama and French, 2000).
-One of the most prominent tasks in this domain is stock movement prediction (Xie et al., 2023a, 2024b, 2024a), which focuses on forecasting the direction of price changes, i.e.
-whether a stock will rise or fall.
-Accurate predictions of financial metrics, including stock movements, interest rates, and economic indicators, are essential for a wide range of stakeholders, from individual investors to large financial institutions.
-However, the inherent complexity and volatility of financial markets pose significant challenges to forecasting, necessitating the use of advanced methodologies to effectively analyze vast and often noisy datasets.
-Traditional stock movement prediction methods (Feng et al., 2018; Qin et al., 2017) relied on analyzing sequences of stock prices and financial indicators [^1] Financial indicators (Jansen, 2020) are professional market analysis metrics derived from price data, where their values or crossover points signal fluctuations in the stock market, indicating rising or falling trends.
-However, relying on one or two indicators can introduce bias, and if an indicator only provides meaningful information at crossover points, it may fail to offer valuable insights during periods without crossings.
-to identify patterns, but they often failed to capture the semantic connections between these sequences, a task at which large language models (LLMs) excel (Huang and Chang, 2022).
-However, LLM-based methods (Xie et al., 2023b, 2024a, a) have primarily shifted focus to leveraging textual data such as news and tweets, often using only the past several days’ closing prices as reference while overlooking the rich insights embedded in historical time-series data (Bustos and Pomares-Quimbaya, 2020).
-This highlights the need to effectively integrate financial time-series data with LLMs.
-Directly working with such data presents a challenge due to its vast scale, encompassing both the breadth of influencing variables and the depth of historical trends.
-To address this, a retrieval-based approach is necessary to efficiently sift through extensive time-series data, extracting meaningful information to enhance LLMs’ ability to deliver more accurate and robust stock movement predictions.
-Although retrieval-augmented generation (RAG) methods have been applied for various tasks (Joshi et al., 2024; He et al., 2024; Cui et al., 2024), it is challenging to directly apply them to financial time-series retrieval.
-First, most embedding-based retrieval methods (retrievers) (Su et al., 2023; Li and Li, 2023; Xiao et al., 2023; Zhang et al., 2023; Wang et al., 2023) are trained on textual data and struggle to process numeric time-series data, which lacks explicit semantic information.
-Second, distance-based retrieval methods like Dynamic Time Warping (DTW) (Yang et al., 2024) focus on numeric similarities and fail to capture deeper semantic relationships or contextual nuances essential for stock movement prediction.
-These limitations underscore the need for a more tailored retrieval framework designed for financial time-series data.
-[…]
-The framework of FinSeer, illustrating its components and workflow.
-Figure 1. Overview of our time-series RAG framework.
-To address these challenges, we introduce a retrieval-augmented framework for stock movement prediction, featuring the first dedicated retriever for financial time-series forecasting, Financial Time-Series Retriever (FinSeer).
-Unlike previous RAG methods, which rely on pre-trained encoders or distance-based metrics, our framework is the first to train a dedicated retriever tailored for continuous and complex temporal sequences.
-The whole retrieval framework is shown in Figure 1.
-To address the limitations of existing datasets, which often contain only price data, and to better align with professional financial analysis, we construct new datasets by integrating 20 groups of financial indicators.
-These indicators provide critical insights into market behavior that stock prices alone cannot capture [^2] For instance, the overbought area in the KDJ indicator group (Wu and Diao, 2015), where $K>80$, $D>70$, and $J>90$, signals a potential falling trend..
-We begin by selecting high-trade-volume U.S.
-stocks, and the indicators are chosen through domain expert consultation and mutual information analysis.
-These indicators are then serialized and divided into train, valid and test sets.
-By training the retriever on datasets enriched with diverse financial indicators, our retrieval framework incorporates financial domain expertise, enabling it to interpret the implications of trending signals that uncover deeper patterns beyond surface-level trend similarities.
-The testing set, designed to mimic real-world economic analysis, ensures a robust evaluation of model performance in practical financial scenarios.
-To train our retriever to identify semantically related sequences and align the retriever’s priorities with those of the LLM, we introduce a novel retrieval mechanism with three key innovations: a new backbone LLM, a novel method for selecting positive and negative candidates, and a new training objective.
-First, to activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA3.2-1B Instruct) (Grattafiori et al., 2024), creating StockLLM as the backbone model of our retriever.
-By using a smaller LLM, we establish a more challenging experimental setup, ensuring performance improvements are attributable to FinSeer rather than the LLM’s capacity.
-Next, StockLLM identifies time-series segments that enhance the generative process, feedbacks from StockLLM are used to select the most beneficial sequences that could lead the model to make the correct decision as positive candidates and the least beneficial as negative candidates.
-This could bridge the gap between the retriever and the LLM and ensure that retrieved data aligns with the LLM’s forecasting priorities.
-Finally, the training objective, inspired by Zhang et al.
-(2023), ensures the retriever distinguishing historically significant sequences (positive candidates) from noisy sequences (negative candidates) by maximizing the similarity between the query and candidate sequences.
-This prioritizes meaningful patterns while minimizing irrelevant data, outperforming traditional methods that struggle to separate signal from noise in complex financial time series.
-Based on FinSeer, the inference process of our RAG framework consists of three steps: (1) using FinSeer to retrieve relevant sequences from the candidate pool, (2) incorporating these sequences into StockLLM’s input context, and (3) combining this enriched context with the original query to perform forecasting.
-This approach effectively integrates salient historical patterns with current data, improving prediction accuracy.
-To thoroughly evaluate the performance of our RAG framework and the retriever, we rely solely on temporal information, excluding additional textual data like news.
-For the RAG framework evaluation, StockLLM serves as the backbone model, integrated with different retrieval methods for comparison.
-Experimental results demonstrate that using our RAG framework outperforms bare StockLLM, indicating that the RAG framework enhances LLM performance by incorporating relevant time-series data as financial context.
-However, RAG with retrieval models trained on text data does not always improve performance compared to bare StockLLM.
-In some cases, performance even declines.
-This suggests that these retrieval models cannot effectively identify beneficial time-series data that could enhance the model’s predictions, which highlights a fundamental gap between text-based retrieval and time-series data.
-Furthermore, RAG framework with FinSeer consistently surpasses all other retrieval models, such as Instructor (Su et al., 2023), BGE (BAAI General Embedding) (Xiao et al., 2023), LLM-Embedder (Zhang et al., 2023), UAE (Li and Li, 2023), and E5-mistral-7b-instruct (Wang et al., 2023), demonstrating the effectiveness of our trained FinSeer in identifying relevant time-series financial data to enhance stock movement prediction.
-Further analysis of retrieval results reveals that FinSeer uniquely identifies relevant financial indicators that enhance LLM performance, while other retrievers predominantly select sequences based on superficial similarities, such as close price and adjusted close price.
-This behavior underscores a critical limitation of text-trained models: they focus on surface-level patterns and fail to discern truly valuable time-series information.
-In contrast, FinSeer effectively bridges this gap, demonstrating its ability to uncover underlying relevance in financial data.
-Our contributions are summarized as follows:
-  1. (1)
-    We propose the first RAG framework for financial time-series forecasting, in which the retrieval mechanism features three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a novel candidate selection method leveraging LLM feedback to identify beneficial sequences, and a training objective that maximizes similarity between queries and historically significant sequences, enabling the retriever to uncover meaningful patterns while minimizing noise in complex financial data.
-  2. (2)
-    We construct new training and testing datasets to address the limitations of existing ones containing only price data, and better align with real-world economic analysis.
-Besides price data, these datasets integrate financial indicators that signal upward or downward trends, carefully selected through domain expert consultation and mutual information analysis.
-By using new training datasets, we train the retriever to uncover deeper patterns beyond surface-level trends during training.
-While the testing set, designed to mimic real-world scenarios, ensures robust evaluation of model performance in practical financial applications.
-  3. (3)
-    We propose a new retriever, FinSeer, specifically designed for time-series forecasting, trained on newly constructed datasets that integrate financial indicators alongside historical stock prices.
-By leveraging LLM feedback and prioritizing historically significant sequences, FinSeer uncovers deeper patterns beyond surface-level trends, setting it apart from traditional methods.
-  4. (4)
-    Experimental results demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, highlighting its effectiveness, while our retriever, FinSeer, surpasses existing RAG methods in stock movement prediction across three datasets, achieving an 8% higher accuracy on BIGDATA22 compared to a general-purpose LLM-feedback-based retriever.
-Furthermore, FinSeer retrieves more dynamic and impactful sequences, showcasing its ability to identify and leverage the most relevant patterns for forecasting.
-2. Problem Definition and Goals
+Retrieval-augmented Large Language Models for Financial Time Series Forecasting
+Abstract
+Accurately forecasting stock price movements is critical for informed financial decision-making, supporting applications ranging from algorithmic trading to risk management.
+However, this task remains challenging due to the difficulty of retrieving subtle yet high-impact patterns from noisy financial time-series data, where conventional retrieval methods, whether based on generic language models or simplistic numeric similarity, often fail to capture the intricate temporal dependencies and context-specific signals essential for precise market prediction.
+To bridge this gap, we introduce FinSrag [^2] Code and data are available at FinSrag ., the first retrieval-augmented generation (RAG) framework with a novel domain-specific retriever FinSeer for financial time-series forecasting.
+FinSeer leverages a candidate selection mechanism refined by LLM feedback and a similarity-driven training objective to align queries with historically influential sequences while filtering out financial noise.
+Such training enables FinSeer to identify the most relevant time-series data segments for downstream forecasting tasks, unlike embedding or distance-based retrieval methods used in existing RAG frameworks.
+The retrieved patterns are then fed into StockLLM, a 1B-parameter LLM fine-tuned for stock movement prediction, which serves as the generative backbone.
+Beyond the retrieval method, we enrich the retrieval corpus by curating new datasets that integrate a broader set of financial indicators, capturing previously overlooked market dynamics.
+Experiments demonstrate that FinSeer outperforms existing textual retrievers and traditional distance-based retrieval approaches in enhancing the prediction accuracy of StockLLM, underscoring the importance of domain-specific retrieval frameworks in handling the complexity of financial time-series data.
+1 Introduction
+Financial time-series forecasting is crucial for maintaining market stability and efficiency, with profound implications for investment decisions, risk assessment, and economic policy formulation (Fama and French, 2000).
+However, the extreme volatility and nonlinear dynamics of financial markets pose significant analytical challenges, requiring sophisticated approaches to distill actionable signals from complex, noise-laden data streams.
+Stock movement prediction, which involves determining future price direction (up or down), represents a particularly critical yet demanding task that has attracted substantial research interest (Xie et al., 2023a, 2024a, 2024b).
+Although recent LLM-based approaches have shown promise in stock prediction, they predominantly rely on textual data like news and social media (Wang et al., 2024), using only the past several days’ closing prices as reference while largely ignoring the wealth of information contained in historical time-series patterns (Bustos and Pomares-Quimbaya, 2020).
+This oversight creates a critical gap in effectively harnessing temporal financial data for prediction.
+The challenge is further compounded by the enormous scale and complexity of such data, spanning multiple influencing factors and extended historical contexts.
+Addressing this requires an intelligent retrieval mechanism capable of efficiently navigating vast time-series datasets to extract and deliver the most relevant patterns, thereby empowering LLMs to generate more accurate and reliable stock movement forecasts.
+Despite the critical importance of retrieving relevant historical patterns for stock movement prediction, current embedding-based and distance-based retrieval-augmented generation (RAG) approaches (Joshi et al., 2024; He et al., 2024; Cui et al., 2024) face fundamental limitations when applied to financial time-series data.
+Conventional embedding-based methods struggle because numeric financial sequences often exhibit superficially similar patterns that lack explicit semantic meaning, making it difficult for text-trained retrievers to distinguish meaningful signals from spurious correlations.
+Meanwhile, distance-based techniques like Dynamic Time Warping (DTW) (Yang et al., 2024) prove inadequate as they typically rely on single-variable comparisons (e.g., adjusted closing prices), ignoring the rich contextual information from other financial indicators.
+This approach limits retrieval to simplistic trend matching and fails to capture deeper, context-dependent relationships where sequences with divergent or opposing trends may actually contain complementary predictive signals.
+The inherent complexity of financial markets demands retrieval methods that can move beyond surface-level pattern matching to identify truly informative historical sequences that may exhibit complex, non-obvious relationships to current market conditions.
+Figure 1: Overview of our time-series RAG framework.
+To address these challenges, we introduce FinSrag, the first Financial time-Series RAG framework for stock movement prediction, using FinSeer, a Financial Time-Series Retriever to effectively retrieve the most beneficial historical sequences beyond surface-level pattern matching for downstream forecasting.
+FinSeer is trained to identify sequences that are embedded with essential or complementary information with limited similarity compared with the query that are beneficial for prediction.
+A key challenge in training FinSeer lies in the absence of explicit retrieval ground truth, that for any given query, it is inherently ambiguous which historical patterns will most effectively aid forecasting given the vast amount of data.
+To address this, we propose an LLM-guided relevance estimation, where the language model itself is used to assess and refine the selection of candidate sequences during training.
+Building on the approach of Zhang et al.
+(2023), we assess candidate relevance by feeding query-candidate pairs into the LLM and using the generation logits corresponding to the correct answer as a proxy for relevance.
+Intuitively, candidates yielding higher logits are considered more informative for the forecasting task.
+We then train FinSeer to effectively distinguish high-value sequences (positives) from irrelevant sequences (negatives) by learning a retrieval embedding space where relevant candidates are closely aligned with their corresponding queries.
+Specifically, for each query, FinSeer is optimized to maximize the similarity score between the query embedding and the embedding of top-ranked candidates (i.e., those associated with higher generation logits from the LLM), while simultaneously pushing apart embeddings of low-ranked, uninformative candidates.
+This contrastive learning objective encourages FinSeer to internalize the LLM’s implicit relevance signals, thereby enhancing its ability to retrieve sequences that are truly beneficial for downstream forecasting.
+Beyond improving relevant sequence retrieval, our framework introduces a crucial advancement in what to retrieve by expanding the candidate pool beyond traditional price-based segments.
+Specifically, we segment the financial time series into candidates, where each candidate corresponds to a time-series segment of a single feature, such as the adjusted close price or a specific technical indicator.
+Unlike prior approaches that retrieve only from price sequences, for the first time, our method incorporates 28 additional financial indicators, allowing retrieval over a richer set of candidate sequences that capture diverse aspects of market behavior [^3] For instance, candidates derived from the KDJ indicator group, particularly in the overbought region where $K>80$, $D>70$, and $J>90$ can reveal early signals of potential reversals (Wu and Diao, 2015)..
+To support this, we construct a new dataset and enriched existing two where time-series segments from both price data and a broad set of financial indicators are included as retrieval candidates, enabling FinSeer to discover deeper predictive patterns beyond surface-level price trends.
+Comprehensive evaluations across multiple financial time-series datasets demonstrate that FinSeer, our specialized retriever, is the only approach that consistently enhances LLM forecasting accuracy, outperforming five state-of-the-art retrievers from the MTEB leaderboard when integrated into the FinSrag framework.
+This superior performance not only validates our domain-specific design for time-series data, but also crucially reveals FinSeer’s unique capability to identify and retrieve the most predictive financial indicator segments.
+Notably, the distance-based method which cannot leverage FinSrag’s architecture, shows weaker performance, further confirming the advantages of our framework’s structured approach to sequence selection, by incorporating diverse financial indicators as candidates.
+More than just improving accuracy, FinSeer’s retrieval patterns provide valuable financial insights: By automatically discovering and leveraging latent market factors that conventional approaches miss, our method offers an empirically grounded alternative to theoretical assumptions in economic models.
+This data-driven approach enables systematic identification of predictive market features, opening new avenues to study market inefficiencies and dynamic dependencies through the lens of actual predictive relevance rather than ex-ante hypotheses.
+In conclusion, our contributions are summarized as follows:
+  1. 1.
+    We introduce FinSrag, the first RAG framework with domain-specific retriever for financial time-series forecasting.
+By improving the capability of identifying and retrieving the most relevant time-series data segments, FinSrag significantly enhances LLMs’ forecasting capabilities in financial markets.
+  2. 2.
+    We propose FinSeer, a novel domain-specific retriever trained via LLM feedback to uncover overlooked market signals.
+We complement this with an enriched dataset featuring 28 expert-selected technical indicators that go beyond conventional price data.
+  3. 3.
+    Experimental results demonstrate the effectiveness of our framework, which surpasseses state-of-the-art retrievers and distance-based methods.
+2 Prelimilaries
 Retrieval-augmented financial time-series forecasting (Sezer et al., 2020) involves predicting future values or trends ($G$) based on a given query sequence ($q$) and a set of retrieved historical sequences ($c$).
@@
 By providing relevant context, the retrieval model enhances the forecasting model’s ability to make accurate and reliable predictions.
-Stock Movement Prediction Task.
 In the specific task of stock movement prediction (Xu and Cohen, 2018), the problem is framed as a binary classification task: predicting whether a stock’s price will rise or fall on the next trading day.
 Given a query sequence $q$, which represents the stock’s price over the previous $t$ days, the model retrieves relevant sequences as context and predicts the stock’s movement $M_{q,d}$ for the next trading day $d$.
-Table 1 defines the major symbols used in this paper.
-Table 1. The definition of symbols.
-General Time-series Symbol Stock Movement Prediction Symbol Definitions
-$q$ $q=\{q_{d-t},...,q_{d-1}\}$ The query time-series data.
-In stock movement prediction, $q$ refers to the query stock price sequence of length $t$, containing stock price data from trading day $d-t$ to trading day $d-1$.
-$G(q)$ $G(q,d)\in\{\textit{rise},\textit{fall}\}$ The final output $G$ given the query $q$.
-In stock movement prediction, $G(q,d)$ shows the generation $G$ of the query stock $q$ on the query trading day $d$, belonging to rise or fall.
-$P(c)=LLM(O|q,c)$ $P(c)=LLM(M_{d}|q,c)$ The probability $P$ of the LLM to generate an accurate output $O$ given the query sequence $q$ and a candidate sequence $c$.
-In stock movement prediction, $P(c)$ refers to generating the accurate movement $M$ on the query trading day $d$.
-=$\mathbb{C_{P}}=\{c_{i}\mid i=1,\ldots,k\}$ The set of top-k retrieved sequences as positive examples, where $P(c_{i})\geq P(c_{i+1})$.
-=$\mathbb{C_{N}}=\{c_{i}\mid i=k+1,\ldots,n\}$ The set of negative retrieved sequences, where $P(c_{i})\geq P(c_{i+1})$.
-$w_{i}$ The soft weight of the $i^{th}$ retrieved sequences, where $w_{i}=P(c_{i}),i=1,...,k$.
-$R$ The retrieve model.
-Rise/Fall Threshold Settings. To classify daily movements as rise or fall, we first calculate returns $R_{t}$, which represents the percentage change in the closing price over consecutive days.
-[…]
-Following Yoo et al.
-(2021) and  Soun et al.
-(2022), we classify the movement as rise if return exceeds 0.55, fall if it is below -0.5.
-In line with previous experimental settings (Xie et al., 2023b, 2024a), we do not evaluate freeze cases.
-However, we include sequences with $R_{t}\in[-0.5,0.55]$ as freeze candidates in the candidate pool, ensuring a diverse and comprehensive set of historical data for context.
-$M_{q,d}&=\begin{cases}\text{rise},&R_{t}>0.55\\ \text{fall},&R_{t}<-0.5\\ \text{freeze},&-0.5\leq R_{t}\leq 0.55\end{cases}\end{aligned}$
-Rationale for Unbalanced Thresholds.
-The asymmetrical thresholds for classifying daily movements (0.55 for rises and -0.5 for falls) reflect the inherent dynamics of stock market behavior.
-Stock prices typically rise gradually due to sustained investor optimism but fall sharply during panic selling or profit-taking.
-The stricter rise threshold prevents minor upward fluctuations from being misclassified as significant increases, while the more lenient fall threshold ensures meaningful downward trends are captured.
-This approach aligns with market realities, improving the reliability of movement classifications.
-3. The RAG Framework
-To identify valuable sequences from vast amounts of stock data and conduct stock movement prediction, we propose a novel RAG framework with new datasets and a novel retrieval mechanism with a new backbone LLM.
-First, we construct three datasets that integrate financial indicators alongside historical stock prices, addressing the limitations of existing datasets and better aligning with real-world economic analysis.
-Second, we introduce StockLLM, a fine-tuned 1B parameter LLM, as the backbone to ensure instruction-following capabilities and establish a challenging experimental setup.
-Finally, we design a novel retrieval mechanism that leverages LLM feedback to select positive and negative candidates and employs a new training objective to prioritize historically significant sequences.
-This comprehensive framework bridges the gap between retrieval models and LLMs, enhancing the time-series forecasting process.
-3.1. Dataset Construction
-To address the limitations of existing datasets and better align with real-world economic analysis, we construct new datasets by integrating financial indicators alongside historical stock prices, enabling deeper insights into market behavior that stock prices alone cannot capture.
-This section details stock selection, raw data collection (Collection of Basic Price Indicators) and financial indicator selection process(Calculation and Selection of Financial Indicators).
-Specifically, our datasets are built using high-trade-volume U.S.
-stocks, while the indicators are selected through domain expert consultation and mutual information analysis.
-Stock Selection.
-To construct our datasets, we select high-trade-volume U.S.
-stocks across three periods: 2014-2015, 2017-2018, and 2022-2023.
-The first two periods align with two benchmark datasets: for 2014-2015, we use the same stocks as the ACL18 dataset (Xu and Cohen, 2018), and for 2017-2018, we use the same stocks as the BIGDATA22 dataset (Soun et al., 2022).
-To incorporate more recent data, we manually select high-trade-volume stocks from 2022-2023 to create the STOCK23 dataset.
-This ensures our training and evaluation reflect recent market conditions and provide a robust benchmark for stock movement prediction.
-Collection of Basic Price Indicators.
-Stock data is collected using the Yahoo Finance API (Xu and Berkely, 2014), which provides basic price indicators such as the opening price, highest price, lowest price, adjusted closing price, and trading volume for each trading day.
-The daily price movement is then calculated as the reference answer for model training and evaluation.
-We then divide the stock data into training, validation, and test sets.
-To ensure temporal integrity and prevent information leakage, we partition the dataset by stocks rather than by time.
-This approach maintains consistent time ranges across splits while ensuring that stocks in the validation and test sets are entirely unseen during training.
-Unlike time-based splits, stock-based partitioning provides a more robust evaluation of the model’s ability to generalize to new stocks, better reflecting real-world scenarios where models are applied to unseen entities.
-Table 2 summarizes the stock counts for each split in our datasets.
-Table 2. Dataset statistics.
-Dataset Stock amount Trading dates
-train valid test all sequences query sequences
-ACL18 (Xu and Cohen, 2018) 33 5 33 2014.06.02-2015.12.31 2015.06.03-2015.12.31
-BIGDATA22 (Soun et al., 2022) 22 3 22 2019.04.01-2020.12.31 2020.04.09-2020.12.31
-STOCK23 24 3 24 2022.01.03-2023.12.31 2023.01.03-2023.12.31
-Calculation and Selection of Financial Indicators.
-In our datasets, we integrate financial indicators to provide critical insights into market behavior that stock prices alone cannot capture.
-Following analyzing methods in the book Machine learning for trading (Jansen, 2020), we conduct an analysis using query sequences in all training sets to explore the relationship between commonly used financial indicators and returns.
-To measure the relationship between these indicators and returns, we utilize Mutual Information (MI), a non-linear measure of dependency that quantifies the amount of information one variable contains about another.
-Specifically, the MI between an indicator $X$ and returns $R_{t}$ is calculated as:
-[…]
-where $p(x,r)$ represents joint probability density function of $X$ (indicator) and $R_{t}$ (returns), $p(x)$ shows marginal probability density function of $X$, and $p(r)$ is marginal probability density function of $R_{t}$.
-We compute MI scores for each indicator and select the top-20 indicator groups with the highest scores as our candidates, with the detailed score ranking provided in Appendix B.1.
-Additionally, an example of a data piece from our datastore is illustrated in Appendix A.1.
-3.2. Sequence Serialization
+Details about threshold settings are illustrated in Appendix C.
+3 Methodology
+3.1 Retrieval Candidate Pool Design and Dataset Construction
+To support our retrieval framework, we introduce a new dataset comprising high-volume stocks from 2022-2023 and enhance two existing datasets with 28 carefully selected financial indicators.
+Complete dataset statistics are presented in Table 5, while partitioning details and additional specifications appear in Appendix E.1.
+The following parts detail our methodology for: (1) stock price data collection and indicator selection (in Collection of Indicators), and (2) temporal scope determination and content specification for queries and candidate sequences (in Query and Candidate Construction).
+Table 1: Dataset statistics of query and candidate sequence amounts [^5] In the table, the candidate rows represent the complete temporal range of each dataset.
+For any specific query, the corresponding candidate pool spans from the dataset’s start date through the trading day immediately preceding the query date, as the example shown in Figure 2(b)..
+Dataset License Trading Dates Train Valid Test
+ACL18 (Xu and Cohen, 2018) MIT License query 2015.06.03-2015.12.31 3,312 440 2,912
+candidate 2014.06.02-2015.12.31 441,694 67,320 444,312
+BIGDATA22 (Soun et al., 2022) Public query 2020.04.09-2020.12.31 3,229 434 2,868
+candidate 2019.04.01-2020.12.31 328,372 44,778 328,372
+STOCK23 MIT License query 2023.01.03-2023.12.31 4,268 570 4,128
+candidate 2022.01.03-2023.12.31 404,736 50,592 404,736
+Collection of Indicators.
+Our financial data foundation comes from the Yahoo Finance API (Xu and Berkely, 2014), which provides 6 daily trading metrics, including opening, highest, lowest, close prices, adjusted close prices, along with trading volume.
+While these basic price data offer essential market snapshots, they fail to capture deeper market dynamics crucial for forecasting.
+For example, if the adjusted close price increases on continuous trading dates, it only shows upward momentum but cannot determine when a trend is exhausted.
+The KDJ indicator, however, combines price momentum and range position to signal overbought (>80) or oversold (<20) conditions Wu and Diao (2015).
+These limitations motivate our multi-layer dataset design to extract deeper and latent market signals beyond raw price movements.
+This multi-layer feature design is inspired by established practices in quantitative finance and machine learning  Heaton et al.
+(2017); Jansen (2020); Chen et al.
+(2024), which emphasize the importance of hierarchical signal extraction from raw market data to enhance model predictive power.
+The dataset consists of three hierarchical components: (1) Core Price Metrics: We begin with aforementioned 6 fundamental indicators, including raw OHLCV (Open, High, Low, Close, Volume) and Adjusted Close prices.
+(2) Primary Technical Indicators: To capture mid-level market dynamics (e.g., trend strength, volatility, and price-volume interactions) Lo et al.
+(2000); Park and Irwin (2007); Wu and Diao (2015), which are not directly observable in raw OHLCV, we incorporate 10 widely-used technical indicators such as log returns, price momentum, and VWAP, whose practice is supported by both empirical studies and industry applications Chan (2013); Fischer and Krauss (2018); Gu et al.
+(2020a); Ma and Yan (2022).
+(3) Alpha Factors: To uncover higher-order predictive signals, we identify 18 Alpha Factors via a Mutual Information (MI)-based selection process Guyon and Elisseeff (2003); Jansen (2020).
+This approach quantifies the nonlinear dependency between candidate features and future returns, allowing us to retain only those indicators with strong predictive relationships.
+The effectiveness of alpha factors in enhancing financial model performance has been extensively supported in the literature Harvey et al.
+(2016); Tulchinsky (2019); Gu et al.
+(2020b); Kakushadze (2016).
+The integrated dataset, exemplified in Appendix E.3, combines these three layers (6 price metrics + 10 technical indicators + 18 alpha factors) to provide a multidimensional view of market conditions.
+Complete specifications for all 34 indicators appear in Table 4.
+Query and Candidate Construction.
+The query construction process involves temporal range determination and query content specification.
+For temporal boundaries, we establish that each query must have sufficient candidate sequences available for retrieval by requiring that the query date occurs at least one year after the start date of the corresponding dataset split.
+This temporal buffer ensures adequate historical context for meaningful retrieval.
+We implement a one-day sliding window across trading days to generate consecutive queries.
+For the example shown in Figure 2(a), in the ACL18 dataset spanning from 2014-06-02 to 2015-12-31, we define queries as those occurring between 2015-06-03 and 2015-12-31.
+This configuration guarantees that even the earliest query (2015-06-03) can access a full year of preceding candidate sequences.
+Each query consists of the stock identifier, query date, and recent market data for query-candidate matching.
+The market data includes adjusted closing prices from the five most recent trading days.
+The five-day window aligns with standard financial practices, where multiples of five are commonly used for price change calculations.
+The candidate construction follows similar temporal and content specifications as queries.
+Candidates are drawn from all available historical data preceding each query date, spanning from the dataset’s start date to the trading day before the query.
+We apply a one-day sliding window to generate sequential candidates, with all stocks in the dataset eligible as candidates regardless of query stock matching.
+Each candidate consists of the stock ticker, price movement direction on a specific trading date, a relevant financial indicator, and corresponding indicator values calculated from the preceding five consecutive trading days.
+Figure 2(b) demonstrates this construction through examples of both rising and falling candidates retrieved from the pool.
+Our approach enables seamless market monitoring through incremental updates to the candidate pool.
+When advancing the query date (e.g., from 2015-12-30 to 2015-12-31), we simply append the new trading day’s data (2015-12-30) to the existing pool (2014-06-02 through 2015-12-29), maintaining all historical candidates without requiring model retraining.
+This efficient update mechanism ensures continuous operation while preserving the complete candidate history.
+Figure 2: Illustration of query and candidate construction.
+(a) illustrates how to construct all queries for a given dataset.
+(b) illustrates the corresponding candidate pool for a given query, and the update of candidate pool with trading date updates.
+3.1.1 Sequence Serialization
 Since stock movement prediction depends on the changes in related features rather than their exact values, we serialize stock prices and financial indicators into a time-series format.
 We use JSON to represent these sequences, as it has been demonstrated to effectively support LLMs in interpreting time-series data (Fang et al., 2024; Singha et al., 2023; Yin et al., 2023).
-Query Sequence Serialization We select query sequences where the query date is at least one year after the start date of the corresponding split in the dataset.
-This ensures that each query has enough candidate sequences for retrieval.
-For example, in the ACL18 dataset, which spans from 2014-06-02 to 2015-12-31, the query sequences are defined as those from 2015-06-03 to 2015-12-31.
-Each query sequence contains the adjusted close price from the previous five trading days along with basic stock information.
-The indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days.
-For example, when inquerying about stock MO on 2015-06-02, the query sequences contains stock name, query date, last five trading dates and their corresponding adjusted close prices.
+The following are two examples of a query and a candidate sequence.
+When inquerying about stock MO on 2015-06-02, the query sequence contains the stock identifier (MO), query date (2015-06-02), and the adjusted close prices of last five trading dates.
 The serialized sequence is shown below:
@@
 "adjusted_close_list": [29.669, 29.9872, 29.8657, 29.6227, 29.6227]}
-Candidate Sequence Serialization As new information becomes available, the candidate pool dynamically incorporates sequences from the most recent trading days.
-For instance, when the query date is 2015-06-03, the candidate pool includes sequences from 2014-06-02 to 2015-06-02.
-When the query date advances to 2015-06-04, the sequence for 2015-06-03 is added to the candidate pool without the need for additional training.
-This demonstrates the scalability of our method.
-For each candidate stock on a specific date, we select 6 basic price indicators and 20 groups of financial indicators.
-Each candidate sequence includes a typical financial indicator along with basic stock information.
-As the same as query serialization, the indicator is represented as a five-day sequence, with a one-day sliding window applied across trading days.
-For example, a candidate represent stock MO on date 2014-07-02, with highest price as its indicator.
-The sequence includes basic information containing stock name, date, movement, last five trading dates and their corresponding highest prices.
+A potential candidate in the candidate pool represent stock MO on date 2014-07-02, with the highest price as its indicator.
+The sequence includes the stock ticker (MO), price movement direction (freeze) on a specific trading date (2014-07-02), a key financial indicator (the highest price), and corresponding indicator values calculated from the preceding five consecutive trading days.
 The serialized sequence is shown below:
@@
 "high_list": [42.2, 42.0, 41.86, 42.28, 42.0]}
-3.3. Retriever Training
-To identify valuable information from vast amounts of sequences and uncover semantic relationships between them, we introduce a novel retrieval mechanism featuring three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a candidate selection method leveraging LLM feedback to identify beneficial sequences and a training objective that maximizes similarity between queries and historically significant sequences, enabling the trained retriever to uncover meaningful patterns while minimizing noise in complex financial data.
-While inspired by (Zhang et al., 2023), our approach is uniquely tailored to financial time-series data, addressing the challenges of relevance assessment and alignment in a domain where traditional retrieval methods often fall short.
-This section details the training process of our framework, including the training of the LLM backbone, candidate scoring and selection, and the training of the retriever, which contains the training objective and knowledge distillation.
-The LLM Backbone.
-To activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA 3.2-1B-Instruct) (Grattafiori et al., 2024) using the LoRA technique for efficient low-rank adaptation, resulting in StockLLM.
-By intentionally using a smaller backbone model, we establish a more challenging experimental setup, ensuring that performance improvements are attributable to FinSeer’s retrieval capabilities rather than the LLM’s capacity.
-The fine-tuning process is implemented using the LlamaFactory framework (Zheng et al., 2024), with the following configuration: a learning rate of 5e-5, a cosine scheduler, gradient accumulation over 8 steps, mixed-precision (fp16) training, and 5 epochs of training with regular evaluation to log metrics and select the best-performing checkpoint.
-This setup demonstrates the robustness of our framework in extracting meaningful insights even under constrained model size, highlighting the effectiveness of FinSeer in enhancing financial forecasting tasks.
-Candidate Scoring.
+3.2 Retriever Training
+We then train FinSeer to effectively distinguish high-value sequences (positives) from irrelevant sequences (negatives) by learning a retrieval embedding space where relevant candidates are closely aligned with their corresponding queries.
+To achieve this, we score and select positive and negative candidates as the retriever training corpus (in Candidate Scoring and Selection), then specify the training objective (in Training Objective) and conduct knowledge distillation (in Knowledge Distillation).
+Candidate Scoring and Selection.
 To determine whether a candidate sequence assists in predicting the movement of the query, we use LLM feedback to score each candidate.
-Specifically, for a given query $q$, we integrate the query sequence and each candidate sequence $c_{i}$ from the candidate pool as concurrent inputs to the LLM.
-The LLM outputs logits, which are unnormalized scores representing the model’s confidence for each possible class (e.g., ”rise” or ”fall”).
+The details of our LLM backbone, StockLLM-1B-Instruct (hereafter referred to as StockLLM), are shown in Appendix F.1.
+Specifically, for a given query $q$, we integrate the query sequence and each candidate sequence $c_{i}$ from the candidate pool as concurrent inputs to the StockLLM.
+Then StockLLM outputs logits, which are unnormalized scores representing the model’s confidence for each possible class (e.g., rise or fall).
 These logits are transformed into probabilities $P(c)$ using the softmax function:
-(4) $P(c)=\frac{e^{z_{c}}}{\sum_{j}e^{z_{j}}},$
-where $z_{c}$ is the logit for the correct class (e.g., ”rise” if the true movement is upward) and $z_{j}$ represents the logits for all possible classes.
+$P(c)=\frac{e^{z_{c}}}{\sum_{j}e^{z_{j}}},$ (1)
+where $z_{c}$ is the logit for the correct class (e.g., rise if the true movement is upward) and $z_{j}$ represents the logits for all possible classes.
 The resulting probability $P(c)$ serves as the score for the candidate $c_{i}$ with respect to the query $q$.
-Candidate Selection.
 We rank the candidate sequences in descending order based on their scores $P(c)$.
@@
 Mathematically, the retriever’s objective is formulated as:
-(5) $R(q)={\arg\max}_{s\in\mathbb{C_{P}}\cup\mathbb{C_{N}}}sup(q,s).$
+$R(q)={\arg\max}_{s\in\mathbb{C_{P}}\cup\mathbb{C_{N}}}sup(q,s).$ (2)
 By focusing on sequences that maximize $sup(q,s)$, the retriever ensures that the most informative and contextually relevant historical sequences are identified.
 Knowledge Distillation.
-To leverage the scoring derived from the LLM, we employ knowledge distillation, which transfers knowledge from the teacher model (LLM) to the student model (retriever) by mimicking the teacher’s output distribution.
-This approach effectively captures nuanced patterns and predictions from the LLM.
+To leverage the scoring derived from StockLLM, we employ knowledge distillation, which transfers knowledge from the teacher model (StockLLM) to the student model (retriever) by mimicking the teacher’s output distribution.
+This approach effectively captures nuanced patterns and predictions from StockLLM.
 Specifically, we minimize the Kullback-Leibler (KL) divergence between the candidate distributions computed using the LLM’s rewards and those predicted by the embedding model.
-For each query $q$ and its candidate list {$\{\mathbb{C_{P}},\mathbb{C_{N}}\}$, we derive the LLM’s rewards for the candidates, denoted as $\{P(c_{i}),i=1,...,n\}$.
+For each query $q$ and its candidate list {$\{\mathbb{C_{P}},\mathbb{C_{N}}\}$, we derive StockLLM’s rewards for the candidates, denoted as $\{P(c_{i}),i=1,...,n\}$.
 To make these rewards suitable for distillation, we normalize them using a softmax function with temperature $\alpha$:
-(6) $w_{i}=\text{softmax}_{R}\left(\frac{P(c_{i})}{\alpha}\right).$
+$w_{i}=\text{softmax}_{R}\left(\frac{P(c_{i})}{\alpha}\right).$ (3)
 The KL divergence is then computed as follows:
-(7) $\min\sum_{c}-w_{i}\times\log\left(\frac{\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{c_{i}}\right\rangle/\tau\right)}{\sum_{{c}^{\prime}\in\mathbb{C}}\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{{c}^{\prime}}\right\rangle/\tau\right)}\right),$
+$\min\sum_{c}-w_{i}\times\log\left(\frac{\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{c_{i}}\right\rangle/\tau\right)}{\sum_{{c}^{\prime}\in\mathbb{C}}\exp\left(\left\langle\boldsymbol{e}_{q},\boldsymbol{e}_{{c}^{\prime}}\right\rangle/\tau\right)}\right),$ (4)
 where $\boldsymbol{e}_{q}$ and $\boldsymbol{e}_{c_{i}}$ are the embeddings of the query $q$ and candidate $c_{i}$, respectively, and $\tau$ is a temperature parameter.
 This loss function optimizes the similarity between the query embedding and the embeddings of the top-ranked candidates, enhancing the retriever’s ability to accurately predict stock price movements.
-3.4. Inference
-The inference process of our RAG framework, built on FinSeer, consists of three key steps to enhance stock movement prediction.
-First, FinSeer retrieves relevant sequences from the candidate pool, leveraging its ability to identify historically significant patterns.
-Second, these retrieved sequences are incorporated into StockLLM’s input context, enriching the model’s understanding of temporal dynamics.
-Finally, the enriched context is combined with the original query to generate forecasts.
-By integrating salient historical patterns with current data, this approach bridges the gap between past trends and present conditions, significantly improving prediction accuracy and robustness in real-world financial scenarios.
-4. Experiment
-4.1. Experimental Settings
+3.3 Inference
+During inference, the key innovation of our FinSrag framework lies in how FinSeer’s retrieval directly enhances StockLLM’s forecasting capability.
+Given a query, FinSeer first identifies the most relevant historical sequences by evaluating both temporal patterns and predictive relationships learned during training, filtering out noisy but numerically similar candidates that typically mislead traditional retrievers.
+These selected sequences are then structured and injected into StockLLM’s context window.
+Crucially, unlike standard RAG that simply concatenates retrieved documents, this end-to-end alignment between retrieval and generation is what enables FinSrag to outperform conventional forecasting pipelines where retrieval and prediction models are optimized separately.
+4 Experiment
+4.1 Experimental Settings
 Datasets.
@@
 These thousand-scale queries help mitigate random bias, ensuring a robust and reliable evaluation of model performance.
+Candidate Pool Settings.
 To ensure a comprehensive evaluation, for each query sequence, we include all sequences containing financial indicators across all stocks in the test set (not limited to the same stock), with data available up to the query date, as potential candidates.
@@
 Baselines.
-To evaluate the efficiency of our RAG framework, we choose bare StockLLM and random retrieval as our baselines.
-To evaluate our retriever FinSeer, we select five competitive RAG models from the top of the MTEB English Leaderboard as baselines: (1) Instructor-large (Su et al., 2023): A 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors.
-(2) UAE-large-v1 (Li and Li, 2023): A 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors.
-(3) BGE-large-en-v1.5 (Xiao et al., 2023): A 335M general embedder pre-trained with RetroMAE (Xiao et al., 2022), encoding sequences into 1024-dimensional tensors.
-(4) LLM-Embedder (Zhang et al., 2023): A 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs.
+To evaluate the efficiency of the FinSrag framework, the bare StockLLM-1B-Instruct without retrieval serves as our baseline to figure out whether the retrieval step enhances StockLLM’s prediction ability.
+To evaluate our retriever FinSeer, we tested other retrieval methods, including random retrieval, DTW distance, and five competitive retrieving models from the top of the MTEB English Leaderboard as baselines, containing: (1) Instructor-large (Su et al., 2023), a 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors.
+(2) UAE-large-v1 (Li and Li, 2023), a 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors.
+(3) E5-mistral-7b-instruct (Wang et al., 2023), a 7111M embedder initialized from Mistral-7B-v0.1 (Jiang et al., 2023) and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
+(4) BGE-large-en-v1.5 (Xiao et al., 2023), a 335M general embedder pre-trained with RetroMAE (Xiao et al., 2022), encoding sequences into 1024-dimensional tensors.
+(5) LLM-Embedder (Zhang et al., 2023), a 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs.
 It encodes sequences into 768-dimensional tensors.
-(5) E5-mistral-7b-instruct (Wang et al., 2023): A 7111M embedder initialized from Mistral-7B-v0.1 (Jiang et al., 2023) and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
+More details are shown in Appendix F.2.
 Evaluation Metrics.
@@
 These metrics evaluate the performance of stock movement prediction based on the distribution of positive and negative samples.
-4.2. Main Results
-Table 3 summarizes the performance of stock movement prediction across the three datasets.
-Table 3. Results of stock movement predictions using LLMs and retrieval models. The asterisk (*) indicates the LLM employed while using retrieval models.
+4.2 Main Results
+As shown in Table 2, experimental results demonstrate our framework’s effectiveness by outperforming all baseline retrieval methods in assisting LLM-based stock movement prediction.
+Table 2: Results of stock movement predictions using StockLLM-1B-Instruct and retrieval models.
 Retrieving Methods ACL18 BIGDATA22 STOCK23
-(+ StockLLM) ACC MCC ACC MCC ACC MCC
-w/o Retrieval 0.496 -0.007 0.497 -0.008 0.509 0.021
-Random Retrieve 0.481 -0.036 0.521 0.045 0.496 -0.004
-DTW 0.518 0.045 0.505 0.021 0.492 -0.007
-Instructor 0.496 -0.008 0.497 -0.006 0.505 0.010
-UAE 0.482 -0.036 0.495 -0.008 0.494 -0.009
-E5 0.510 0.028 0.503 0.008 0.499 -0.001
-BGE 0.489 -0.019 0.487 -0.026 0.488 -0.014
-LLM Embedder 0.499 0.000 0.459 -0.083 0.503 0.007
-FinSeer 0.521 0.042 0.538 0.079 0.542 0.085
-We analyze our framework and retriever separately to evaluate their contributions.
-First, our RAG framework outperforms both bare StockLLM and random retrieval across all datasets, demonstrating its effectiveness in enhancing stock movement prediction.
-Second, among all evaluated retrieval methods, FinSeer consistently achieves the best performance when integrated with StockLLM, showcasing its ability to retrieve valuable sequences that improve forecasting accuracy.
-Notably, FinSeer is the only retriever that achieves stable improvements over bare StockLLM on all datasets, while other methods exhibit inconsistent results, sometimes enhancing predictions and other times degrading them.
-Furthermore, our experiments reveal that incorporating a retrieval process does not always guarantee performance improvements, as retrieval errors can mislead the LLM and harm predictions.
-This underscores the critical need for retrieval models that align closely with the LLM’s objectives, a challenge that FinSeer effectively addresses through its tailored design.
-Figure 2. Indicator occurrences of different RAG models on ACL18 dataset.
-The minimal performance difference between the instruction-finetuned retriever (Instructor) and the no-retrieval baselines highlights the inherent challenge of time-series retrieval.
-Unlike text retrieval, time-series retrieval does not solely rely on task understanding, as candidate sequences often appear similar, making it difficult to distinguish valuable ones.
-The angle-optimized retriever (UAE) shows inconsistent performance across datasets, as its ability to differentiate similarity between query and candidate sequences does not fully align with the LLM’s perception of importance.
-Similarly, LLM Embedder, our backbone model trained with LLM feedback, struggles to generalize to time-series retrieval tasks, further underscoring the complexity of the problem.
-The best-performing baseline is E5, which achieves competitive results despite its significantly larger model size and embedding dimensions (five times larger than FinSeer).
-However, FinSeer outperforms E5 on all datasets.
-This demonstrates that FinSeer successfully learns the preferences of LLMs and effectively enhances their time-series forecasting ability through retrieval, even under challenging conditions.
-In summary, while RAG methods do not guarantee performance improvements due to the risk of retrieval errors, FinSeer stands out by consistently retrieving high-quality sequences that align with the LLM’s forecasting objectives, leading to superior results compared to other retrievers.
-4.3. Ablation Study
-Indicator Occurrences In this part, we analyze what indicators all RAG models retrieve. We calculate indicator occurrences on ACL18 dataset and the results are shown in Figure 2.
+(+ StockLLM-1B-Instruct ) ACC MCC ACC MCC ACC MCC
+w/o Retrieval 0.498 -0.006 0.493 -0.017 0.509 0.021
+Random Retrieval 0.485 -0.028 0.495 -0.007 0.496 -0.004
+DTW 0.516 0.041 0.500 0.010 0.492 -0.007
+Instructor (Su et al., 2023) 0.498 -0.005 0.493 -0.015 0.505 0.010
+UAE (Li and Li, 2023) 0.486 -0.029 0.493 -0.011 0.494 -0.009
+E5 (Wang et al., 2023) 0.510 0.027 0.498 -0.002 0.499 -0.001
+BGE (Xiao et al., 2023) 0.492 -0.012 0.501 0.002 0.488 -0.014
+LLM Embedder (Zhang et al., 2023) 0.503 0.007 0.459 -0.083 0.503 0.007
+FinSeer 0.517 0.035 0.510 0.023 0.542 0.085
+First, compared with bare StockLLM and random retrieval, FinSeer demonstrates the critical importance of retrieving truly valuable information.
+While trained on a comprehensive stock movement prediction corpus, bare StockLLM’s limited input with only the recent five-day adjusted close price results in unstable performance that fluctuates around random guessing levels.
+Similarly, when randomly retrieved sequences are provided as supposedly relevant context, they introduce instability by arbitrarily confusing or occasionally coincidentally benefiting StockLLM’s decision-making.
+Second, our comparison with the five top-ranked retrievers from the MTEB English leaderboard reveals FinSeer’s superiority in time-series retrieval.
+The negligible performance gap between instruction-finetuned retriever (Instructor) and no-retrieval baselines underscores the fundamental challenges of time-series retrieval.
+Unlike text retrieval, this task cannot rely solely on task understanding since candidate sequences often exhibit visual similarity while differing in predictive value.
+Other retrievers demonstrate inconsistent cross-dataset performance because their similarity differentiation fails to align with the LLM’s perception of importance.
+Even LLM Embedder, our backbone model trained with LLM feedback, shows limited generalization to time-series retrieval, further emphasizing the problem’s complexity.
+Among those retrievers, FinSeer consistently outperforms other retrievers across all datasets, proving its superior ability to learn LLM preferences and effectively enhance time-series forecasting through retrieval.
+Third, our framework’s advantages over distance-based retrieval methods highlight the value of incorporating diverse feature types.
+While DTW achieves comparable performance to FinSeer during specific market periods (ACL18 [2014-2015] and BIGDATA22 [2019-2020]), its retrieval capability proves insufficient during the volatile 2022-2023 period (STOCK23) marked by significant market surges and fluctuations.
+During this challenging phase, DTW significantly degrades StockLLM’s performance, while FinSeer maintains its enhancement capability.
+4.3 Ablation Study
+In this section, we explore two aspects of our framework based on retrieval results: which indicators are most retrieved by all retrievers (in Indicator Occurrences), and whether StockLLM-1B-Instruct makes predictions by analyzing candidates or just based on the candidates’ movements (in Appendix G.1).
+We also visualize indicator sequence embeddings in Appendix G.2.
+Moreover, we explore the performance of these retrieval methods with a larger size of StockLLM (in Appendix G.3).
+In this section, we analyze the retrieved indicators of all RAG models.
+We calculate indicator occurrences on the ACL18 test set, and the results are shown in Figure 3.
 As shown in the figure, FinSeer is the only model that successfully extracts a diverse and comprehensive set of indicators while achieving superior performance.
@@
 This richer set of retrieved indicators provides FinSeer with more comprehensive auxiliary information, enabling more accurate and reliable predictions.
-Candidate Movement Correlation In this section, we investigate whether the LLM relies on the movement trends of retrieved sequences to make predictions.
+Figure 3: Indicator occurrences of different RAG models on ACL18 dataset.
+4.4 Case Study
+This case study illustrates the critical importance of alignment between the retriever and the LLM’s forecasting preferences in financial time-series analysis.
+We examine the stock XOM on 2015-06-25 from the ACL18 dataset, where the adjusted close price exhibited a pronounced downward trend.
+The query sequence is as follows:
+{ "query_stock": "XOM",
+"query_date": "2015-06-25",
+"recent_date_list": ["2015-06-18", "2015-06-19", "2015-06-22", "2015-06-23", "2015-06-24"],
+"adjusted_close_list": [58.0813, 57.8979, 57.8707, 57.8027, 57.5377]}
+While multiple retrievers were evaluated, only FinSeer successfully enabled StockLLM to predict the correct movement as a fall.
+Specifically, FinSeer retrieved five diverse indicators: close price, adjusted close price, alpha021, alpha054, and the highest price, providing a comprehensive view of the stock’s behavior.
+Alpha021 identifies trends based on short- and long-term price averages and volume conditions, while alpha054 combines price and volume rankings to assess performance within a specific time window.
+These indicators allowed StockLLM to accurately assess whether the downward trend would persist or reverse, demonstrating the value of retrieving contextually relevant and diverse features.
+In contrast, other retrievers, such as Instructor, BGE, LLM Embedder, and E5, extracted sequences dominated by close or adjusted close prices, all reflecting similar downward trends.
+While these sequences aligned with the current trend, they failed to provide actionable insights for forecasting future movements, leading StockLLM to misinterpret them as noise and incorrectly predict a rise.
+Similarly, UAE retrieved sequences indicating overbought and oversold conditions, including three rise and two freeze trends.
+Although overbought signals often suggest a potential downturn, the retrieved sequences themselves exhibited rising or frozen trends, confusing StockLLM and resulting in an erroneous prediction.
+This case study underscores the superiority of FinSeer in retrieving meaningful and diverse indicators that align with the LLM’s forecasting logic, enabling more accurate and reliable predictions.
+5 Conclusion
+In this paper, we present FinSrag, the first retrieval-augmented generation framework tailored for financial time-series forecasting.
+At its core, FinSeer, a novel retriever refined by LLM feedback, effectively identifies historically influential market sequences while filtering out financial noise.
+Combined with StockLLM, a fine-tuned LLM with 1B parameters, our framework leverages enriched financial datasets to capture previously overlooked market dynamics.
+Empirical results confirm that FinSeer surpasses both textual and distance-based retrievers in improving StockLLM’s prediction accuracy, highlighting the necessity of domain-specific retrieval in financial forecasting.
+Beyond stock markets, FinSrag establishes a blueprint for integrating RAG in time-sensitive decision-making domains.
+Appendix A Limitation
+While our FinSrag framework demonstrates significant improvements in LLM-based time-series forecasting, several limitations suggest important directions for future research.
+First, although our current evaluation focuses specifically on stock movement prediction, the core methodology and architectural principles of our framework are inherently generalizable to other time-series prediction tasks.
+Future work will systematically validate this generalizability across domains, including but not limited to economic indicators, energy demand forecasting, and epidemiological trend prediction.
+Second, the present study deliberately focuses on unimodal time-series data to provide clear, controlled evidence for the effectiveness of our retrieval mechanisms.
+This focused approach has allowed us to rigorously quantify the contribution of temporal pattern retrieval independent of other factors.
+However, we recognize that many real-world forecasting scenarios could benefit from multimodal integration, where time-series data interacts with textual reports, visual data, or other information modalities.
+Future work will incorporate cross-modal corpora to enable more comprehensive information retrieval.
+Appendix B Ethical Statement
+Our framework and its contents are provided for academic and educational purposes only.
+None of the material constitutes financial, legal, or investment advice.
+No warranties, express or implied, are offered regarding the accuracy, completeness, or utility of the content.
+The authors and contributors are not responsible for any errors, omissions, or any consequences arising from the use of the information herein.
+Users should exercise their own judgment and consult professionals before making any financial, legal, or investment decisions.
+The use of the software and information contained in this repository is entirely at the user’s own risk.
+FinSrag and all its components are licensed under MIT.
+Appendix C Task Definition Details
+Rise/Fall Threshold Settings. To classify daily movements as rise or fall, we first calculate returns $R_{t}$, which represents the percentage change in the closing price over consecutive days.
+[…]
+Following Yoo et al.
+[2021] and  Soun et al.
+[2022], we classify the movement as rise if return $R_{t}$ exceeds 0.55, fall if it is below -0.5.
+In line with previous experimental settings [Xie et al., 2023b, 2024b], we do not evaluate freeze cases as queries.
+However, we include sequences with $R_{t}\in[-0.5,0.55]$ as freeze candidates in the candidate pool, ensuring a diverse and comprehensive set of historical data for context.
+$M_{q,d}&=\begin{cases}\text{rise},&R_{t}>0.55\\ \text{fall},&R_{t}<-0.5\\ \text{freeze},&-0.5\leq R_{t}\leq 0.55\end{cases}\end{aligned}$
+Rationale for Unbalanced Thresholds.
+The asymmetrical thresholds for classifying daily movements (0.55 for rises and -0.5 for falls) reflect the inherent dynamics of stock market behavior.
+Stock prices typically rise gradually due to sustained investor optimism but fall sharply during panic selling or profit-taking.
+The stricter rise threshold prevents minor upward fluctuations from being misclassified as significant increases, while the more lenient fall threshold ensures meaningful downward trends are captured.
+This approach aligns with market realities, improving the reliability of movement classifications.
+Appendix D Related Work
+D.1 Stock Movement Prediction
+Non-LLM Methods.
+Traditional approaches to stock movement prediction have focused on various aspects of financial data.
+One prominent category of methods analyzes stock price sequences and their corresponding technical indicators [Qin et al., 2017, Feng et al., 2018] to identify patterns in historical data for predicting future movements.
+However, due to the complexity of factors influencing stock prices, subsequent methods have incorporated additional contextual information, such as news articles [Ding et al., 2015, Liu et al., 2018] or social media posts [Xu and Cohen, 2018, Wu et al., 2018].
+Despite efforts, these methods are highly susceptible to noise and struggle to analyze the vast and diverse nature of financial information effectively.
+LLM-based Methods.
+Recent studies have explored using LLMs for financial prediction tasks, either by fine-tuning open-source models or prompting advanced models like GPT-4.
+However, even state-of-the-art models, including GPT-4, have achieved only random-guessing-level accuracy in stock movement prediction [Xie et al., 2023b, 2024b, a].
+This highlights the inherent challenges in identifying and analyzing meaningful patterns in a domain as volatile and multifaceted as stock market prediction.
+D.2 Time-series forecasting with LLMs
+Non-retrieval Methods.
+To enhance the performance of LLMs in time-series forecasting, existing methods have primarily focused on aligning temporal and textual data, either by transforming time-series into textual formats or by encoding both modalities into a unified vector space.
+For instance, a study [Jin et al., 2023] reprograms time-series data into textual representations suitable for LLMs, enhancing prediction accuracy via declarative prompts.
+Similarly, some other studies [Yu et al., 2023, Liu et al., 2024] explore cross-modal alignment: the former applies LLMs to financial forecasting by integrating stock prices with news data, while the latter introduces a cross-modality framework to align time-series with text for improved predictive performance.
+While these advancements improve alignment, existing methods often struggle to process large-scale time-series data comprehensively due to LLM input limitations.
+As a result, effectively leveraging extensive historical data remains a challenge.
+This highlights the need for retrieval-augmented methods, a gap our approach directly addresses.
+RAG Methods.
+With the rapid advancement of RAG techniques in various applications of LLMs, recent research has begun exploring RAG for time-series forecasting.
+For instance, TimeRAG [Yang et al., 2024] integrates RAG into time-series forecasting by combining Dynamic Time Warping (DTW) with LLMs to improve prediction accuracy.
+However, relying solely on numeric similarity is insufficient for financial time-series forecasting, as it fails to capture deeper semantic relationships.
+This underscores the need for a more targeted retrieval framework tailored to the complexities of financial data, a challenge our framework effectively addresses.
+Appendix E Dataset Construction Details
+E.1 Dataset Partition
+To construct our datasets, we select high-trade-volume U.S.
+stocks across three periods: 2014-2015, 2017-2018, and 2022-2023.
+The first two periods align with two benchmark datasets: for 2014-2015, we use the same stocks as the ACL18 dataset [Xu and Cohen, 2018], and for 2017-2018, we use the same stocks as the BIGDATA22 dataset [Soun et al., 2022].
+To incorporate more recent data, we manually retrieve high-trade-volume stocks from 2022 to 2023 to create the STOCK23 dataset.
+This ensures our training and evaluation reflect recent market conditions and provide a robust benchmark for stock movement prediction.
+To maintain temporal integrity and prevent information leakage, we implement stock-wise partitioning rather than temporal splitting of the dataset.
+This approach addresses a critical characteristic of our data where many high-trade-volume stocks (e.g., Google (GOOG)) appear across multiple years (2014-2015 and 2019-2020).
+Table 3 details the complete stock lists and their corresponding sectors for all subsets, demonstrating two key properties of our partitioning: first, the complete absence of overlap between training and test sets, and second, comprehensive sector coverage in both sets.
+The alphabetical partitioning serves as a neutral, systematic approach that is uncorrelated with any financial or temporal characteristics while achieving our core objectives of clean separation and representative sampling.
+This method guarantees that the model encounters truly novel stocks during testing without any temporal leakage, while maintaining balanced sector representation across all partitions.
+Table 3: Stock information in datasets.
+Dataset Split (amount) Stock List Sector Count
+train (33) ABBV, AEP, AMGN, AMZN, BA, BABA, BAC, BCH, BHP, BP, BRK-A, BSAC, BUD, C, CAT, CHTR, CMCSA, CODI, CSCO, CVX, D, DHR, DIS, DUK, EXC, GD, GE, GOOG, HD, HON, HSBC, IEP, INTC basic-materials (1), communication-services (4), consumer-cyclical (3), consumer-defensive (1), energy (3), financial-services (6), healthcare (3), industrials (6), technology (2), utilities (4)
+ACL18 valid (5) JNJ, JPM, KO, LMT, MA consumer-defensive (1), financial-services (2), healthcare (1), industrials (1)
+test (33) MCD, MDT, MMM, MO, MRK, MSFT, NEE, NGG, NVS, ORCL, PCG, PEP, PFE, PG, PM, PPL, REX, SLB, SNY, SO, SPLP, SRE, T, TM, TSM, UL, UNH, UPS, V, VZ, WFC, WMT, XOM basic-materials (1), communication-services (2), consumer-cyclical (2), consumer-defensive (6), energy (2), financial-services (2), healthcare (6), industrials (3), technology (3), utilities (6)
+train (22) AEP, AMGN, AMZN, BA, BAC, C, CAT, CODI, CSCO, CVX, D, DIS, DUK, EXC, GD, GE, GMRE, GOOG, HD, HON, INTC, JNJ communication-services (2), consumer-cyclical (2), energy (1), financial-services (2), healthcare (2) industrials (6), real-estate (1), technology (2), utilities (4)
+BIGDATA22 valid (3) JPM, KO, LMT financial-services (1), consumer-defensive (1), industrials (1)
+test (22) MA, MCD, MDT, MMM, MO, MRK, MSFT, NEE, ORCL, PCG, PM, PPL, REX, SO, SRE, T, UPS, V, VZ, WFC, WMT, XOM basic-materials (1), communication-services (2), consumer-cyclical (1), consumer-defensive (3), energy (1), financial-services (3), healthcare (2), industrials (2), technology (2), utilities (6)
+train (24) ADBE, ADSK, AMD, AVGO, COIN, COST, DELL, DXCM, ELV, ETSY, FOX, HOOD, HPQ, IBM, INTU, JBL, JNJ, KKR, KMX, LDOS, LLY, LOW, LULU, LUV communication-services (1), consumer-cyclical (4), consumer-defensive (1), financial-services (3), healthcare (4), industrials (1), technology (10)
+STOCK23 valid (3) MARA, META, MNST financial-services (1), communication-services (1), consumer-defensive (1)
+test (24) MRVL, MS, MSCI, NDAQ, NKE, NVDA, ORLY, PANW, PEAK, PLTR, PNC, PNW, RY, SMCI, SNOW, SNPS, T, TGT, TSLA, UBER, ULTA, V, WBA, WDAY communication-services (1), consumer-cyclical (4), consumer-defensive (1), financial-services (6), healthcare (2), technology (9), utilities (1)
+E.2 Financial Indicators
+Financial Indicators Selection We compute the MI scores for each indicator using nonparametric estimation methods provided by mutual_info_regression from the Scikit-learn library.
+For each query with retrieved candidates, we convert the movements of retrieved sequences and LLM’s prediction to numerical representations.
+${M_{c_{j}},M_{q}}&=\begin{cases}\text{1},&\text{rise}\\ \text{0},&\text{freeze}\\ \text{-1},&\text{fall}\end{cases}\end{aligned}$
+Then, we calculate the average value of the $M_{c_{i}}$ from the five retrieved sequences to represent the RAG-provided result, and the final prediction of the query is recorded as another numerical value:
+[…]
+Lastly, we adopt the Pearson correlation coefficient to compute the correlation:
+[…]
+Then we normalize them into range $(0,1)$.
+These scores provide insight into the strength of the dependency between each financial indicator and the forward return.
+Higher MI scores indicate stronger relationships, highlighting which indicators are most predictive of future price movements.
+We select the top-18 alpha indicators with the highest MI scores in our candidates, as is shown in Figure 4.
+Figure 4: Mutual Information(MI) scores between indicators and forward return.
+Indicators Descriptions We provided a brief description of each indicator in Table 4.
+Table 4: Brief descriptions of indicators.
+Group Indicator Description
+basic price indicators
+1 open (opening price) CoinAPI [2023] reflecting the stock’s value at the start of the trading session.
+high, low (highest/lowest price) CoinAPI [2023] capture the price range during the day, offering insights into volatility.
+close (close price) CoinAPI [2023] often the most important for technical analysis, represents the final trading price for the day.
+adj_close (adjusted close price) Ganti [2020] accounts for corporate actions like dividends or stock splits, providing a more accurate long-term price history.
+volume CoinAPI [2023] indicating the level of trading activity and movement, typically the change in price from the previous day, reflects short-term price fluctuations.
+devived financial indicators
+2 movement Tardi [2022], Jegadeesh [1990], Jegadeesh and Titman [1993] represents the change in stock price between consecutive trading days.
+returns Tardi [2022], Jegadeesh [1990], Jegadeesh and Titman [1993] represents the percentage change in the closing price over consecutive days.
+VWAP (Volume Weighted Average Price) Mitchell et al. [2020] calculates the average price of a stock weighted by trading volume, providing a benchmark for institutional investors.
+3 macd_histogram (values) Anghel [2015] measures the difference between a stock’s MACD line and its signal line, helping identify trends and momentum.
+macd_crossover (signals) Anghel [2015] indicates a buy or sell signal when the MACD line crosses above or below the signal line.
+4 bollinger_bands (signals) Bollinger [2001] indicates whether the adjusted close price exceeds the upper or lower price limits around a moving average, serving as a measure of price volatility.
+exceeding_upper/lower (values) Bollinger [2001] the value of the adjusted close price exceeds the upper or lower price limits around a moving average.
+5 overbought_and_oversold_conditions Wilder [1978] classify the market state into "overbought" or "oversold" areas based on predefined thresholds for $K$, $D$, and $J$.
+kdj_crossover (signals) Wu and Diao [2015] identify "bullish" or "bearish" signals based on the relationship between $K$ and $D$.
+6 alpha_smr Kakushadze [2016] measures short-term reversal patterns by capturing stocks’ tendency to revert after abrupt price movements.
+alpha_mom Kakushadze [2016] tracks momentum effects by identifying stocks with strong recent performance likely to continue trending.
+alpha_009 Kakushadze [2016] uses recent price changes to classify trends based on their range within a 5-day window.
+alpha_012 Kakushadze [2016] relates volume change direction to the negative change in price.
+alpha_041 Kakushadze [2016] measures the difference between price range midpoint and VWAP.
+alpha_049 Kakushadze [2016] detects strong downtrends or reversals in lagged price movements.
+alpha_051 Kakushadze [2016] quantifies mean-reversion behavior over medium horizons, typically exploiting price corrections after sustained moves.
+alpha_054 Kakushadze [2016] combines momentum and volume signals to detect stocks with improving trends supported by trading activity.
+alpha_101 Kakushadze [2016] normalizes price change by daily price range.
+7 alpha_002 Kakushadze [2016] measures the negative 6-day correlation between ranked changes in logarithmic volume and ranked price movements.
+alpha_006 Kakushadze [2016] computes the negative correlation between opening price and volume over a 10-day window.
+alpha_024 Kakushadze [2016] compares long-term and short-term price movements to identify breakouts or retracements.
+alpha_028 Kakushadze [2016] scales the relationship between average price, low price correlation, and close price.
+alpha_032 Kakushadze [2016] combines deviations from a 7-day moving average and long-term correlation between VWAP and lagged close.
+alpha_053 Kakushadze [2016] identifies stocks exhibiting strong recent performance (typically over 3-6 months) while controlling for volatility, aiming to capture persistent trends while mitigating risk from erratic price swings.
+8 alpha_021 Kakushadze [2016] identifies trends based on short- and long-term price averages and volume conditions.
+alpha_023 Kakushadze [2016] tracks changes in high prices if they exceed their 20-day average.
+alpha_046 Kakushadze [2016] analyzes trends in lagged and current 10-day price changes.
+E.3 Sequences and Serialization
+An example in our datastore is shown below.
+"stock_name":"ABBV",
+"query_date":"2014-06-05",
+"movement":"rise",
+"open":54.549999,
+"high":55.32,
+"low":54.360001,
+"close":55.299999,
+"adj_close":36.980961,
+"volume":4847300,
+"macd_histogram":0.0621141602,
+"macd_crossover":null,
+"bollinger_bands":null,
+"exceeding_upper":null,
+"exceeding_lower":null,
+"overbought_and_oversold_conditions":null,
+"kdj_crossover":null,
+"returns":0.0131913094,
+"VWAP":54.9933333333,
+"alpha_smr":0.000549858,
+"alpha_mom":0.0121661224,
+"alpha_002":null,
+"alpha_006":null,
+"alpha_009":-0.481476,
+"alpha_012":-0.481476,
+"alpha_021":1,
+"alpha_023":0.0,
+"alpha_024":-0.769035,
+"alpha_028":null,
+"alpha_032":null,
+"alpha_041":-0.1554335256,
+"alpha_046":-0.481476,
+"alpha_049":-0.481476,
+"alpha_051":-0.481476,
+"alpha_053":null,
+"alpha_054":126.4264005641,
+"alpha_101":-18.282056485
+Appendix F Experimental Settings
+F.1 StockLLM-1B-Instruct Backbone
+To activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA 3.2-1B-Instruct) [Grattafiori et al., 2024] using the LoRA technique for efficient low-rank adaptation, resulting in StockLLM-1B-Instruct.
+By intentionally using a smaller backbone model, we establish a more challenging experimental setup, ensuring that performance improvements are attributable to FinSeer’s retrieval capabilities rather than the LLM’s capacity.
+More details of fine-tuning are illustrated in Appendix F.1.
+The fine-tuning process is implemented using the LlamaFactory framework [Zheng et al., 2024], with the following configuration: a learning rate of 5e-5, a cosine scheduler, gradient accumulation over 8 steps, mixed-precision (fp16) training, and 5 epochs of training with regular evaluation to log metrics and select the best-performing checkpoint.
+This setup demonstrates the robustness of our framework in extracting meaningful insights even under constrained model size, highlighting the effectiveness of FinSeer in enhancing financial forecasting tasks.
+The fine-tuning prompt is shown in Figure 5, with candidate sequences randomly retrieved from the dataset.
+The fine-tuning step is to train the StockLLM-1B-Instruct to follow instructions, without additional steps for retrieving certain candidates.
+F.2 Experimental Settings
+Baselines. The introduction of retrieval baselines is as follows:
+  • Random retrieval: The candidate range is aligned with the "Candidate Pool Settings" part. We use the random.sample() function in Python to retrieve random candidates from the candidate pool.
+  • DTW (Dynamic Time Warping) distance: The formula measures the similarity between two temporal sequences by aligning them non-linearly in time, minimizing the cumulative warping cost.
+$D_{\text{DTW}}(X,Y)=\min_{\pi\in\mathcal{P}}\sqrt{\sum_{(i,j)\in\pi}(x_{i}-y_{j})^{2}}$, where $X,Y$ are sequences, $\pi$ is a warping path, and $\mathcal{P}$ is the set of all possible paths.
+  • Instructor-large [Su et al., 2023]: A 335M instruction-finetuned text embedder that encodes sequences into 768-dimensional tensors.
+  • UAE-large-v1 [Li and Li, 2023]: A 335M ANGLE-optimized text embedding model that encodes sequences into 1024-dimensional tensors.
+  • E5-mistral-7b-instruct [Wang et al., 2023]: A 7111M embedder initialized from Mistral-7B-v0.1 [Jiang et al., 2023] and fine-tuned on multilingual datasets, encoding sequences into 4096-di-mensional tensors.
+  • BGE-large-en-v1.5 [Xiao et al., 2023]: A 335M general embedder pre-trained with RetroMAE [Xiao et al., 2022], encoding sequences into 1024-dimensional tensors.
+  • LLM-Embedder [Zhang et al., 2023]: A 109M embedder fine-tuned from BGE-large-en-v1.5, designed as a unified embedding model to support diverse retrieval augmentation needs for LLMs.
+It encodes sequences into 768-dimensional tensors.
+Prompts. The prompt for bare StockLLM-1B-Instruct is shown in Figure 6, and the prompt for retrieval methods is shown in Figure 5.
+Figure 5: The prompt template for StockLLM-1B-Instruct fine-tuning and RAG methods evaluation.
+Figure 6: The prompt template for bare StockLLM-1B-Instruct evaluation.
+Hyperparameters.
+For each model, we employ default parameter settings, utilizing official models for open-source LLMs obtained from Hugging Face.
+These testing procedures take place on a computational infrastructure consisting of one NVIDIA A800 Tensor Core GPU, equipped with 80GB of memory.
+Error Bars.
+To account for potential variability, we conducted three randomized retrieval trials and report averaged performance metrics.
+For similarity-based retrieval methods, since both queries and candidates are encoded as fixed vector representations, the retrieved candidates remain deterministic across runs.
+To verify this stability, we performed three identical inference passes through StockLLM, observing minimal variance in prediction outcomes.
+Appendix G Addtional Experimental Results
+G.1 Candidate Movement Correlation
+In this section, we investigate whether StockLLM-1B-Instruct relies on the movement trends of retrieved sequences to make predictions.
 To this end, we compute the correlation between the movements of retrieved sequences and the LLM’s generated results.
-The detailed calculation process is provided in Appendix B.2.
 The results in Table 5 reveal no significant correlation between the movement direction (rise or fall) of the retrieved candidate sequences and the final predictions of StockLLM.
-This finding indicates that StockLLM does not simply mirror the movement trends of the retrieved sequences but instead analyzes their specific content to infer the query’s movement direction.
+This finding indicates that StockLLM-1B-Instruct does not simply mirror the movement trends of the retrieved sequences but instead analyzes their specific content to infer the query’s movement direction.
 This ability highlights the LLM’s capacity to extract meaningful insights from complex time-series data.
-Table 4. Correlation between the movements of retrieved sequences and LLM’s generated results.
+Table 5: Correlation between the movements of retrieved sequences and LLM’s generated results.
 Dataset RAG Model Correlation P-Value
@@
 This observation underscores the critical role of retrieval quality in RAG models, as the relevance and informativeness of retrieved sequences directly influence prediction performance.
-Notably, StockLLM is a relatively small 1B parameter LLM, making this a particularly challenging setting.
+Notably, StockLLM-1B-Instruct is a relatively small 1B parameter LLM, making this a particularly challenging setting.
 Despite this, our time-series-tailored RAG model, FinSeer, successfully enhances StockLLM’s performance by retrieving rich and relevant time-sensitive information.
 This demonstrates the effectiveness of our approach in improving prediction accuracy and showcases its potential for financial time-series forecasting.
-4.4. Case Study
-This case study illustrates the critical importance of alignment between the retriever and the LLM’s forecasting preferences in financial time-series analysis.
-We examine the stock XOM on 2015-06-25 from the ACL18 dataset, where the adjusted close price exhibited a pronounced downward trend.
-The query sequence is as follows:
-{ "query_stock": "XOM",
-"query_date": "2015-06-25",
-"recent_date_list": ["2015-06-18", "2015-06-19", "2015-06-22", "2015-06-23", "2015-06-24"],
-"adjusted_close_list": [58.0813, 57.8979, 57.8707, 57.8027, 57.5377]}
-While multiple retrievers were evaluated, only FinSeer successfully enabled StockLLM to predict the correct movement as a fall.
-Specifically, FinSeer retrieved five diverse indicators—close price, adjusted close price, alpha021, alpha054, and the highest price—providing a comprehensive view of the stock’s behavior.
-Alpha021 identifies trends based on short- and long-term price averages and volume conditions, while alpha054 combines price and volume rankings to assess performance within a specific time window.
-These indicators allowed StockLLM to accurately assess whether the downward trend would persist or reverse, demonstrating the value of retrieving contextually relevant and diverse features.
-In contrast, other retrievers, such as Instructor, BGE, LLM Embedder, and E5, extracted sequences dominated by close or adjusted close prices, all reflecting similar downward trends.
-While these sequences aligned with the current trend, they failed to provide actionable insights for forecasting future movements, leading StockLLM to misinterpret them as noise and incorrectly predict a rise.
-Similarly, UAE retrieved sequences indicating overbought and oversold conditions, including three rise and two freeze trends.
-Although overbought signals often suggest a potential downturn, the retrieved sequences themselves exhibited rising or frozen trends, confusing StockLLM and resulting in an erroneous prediction.
-This case study underscores the superiority of FinSeer in retrieving meaningful and diverse indicators that align with the LLM’s forecasting logic, enabling more accurate and reliable predictions.
-4.5. Indicator Sequence Visualization
-To intuitively examine how FinSeer embeds indicator sequences, we use LargeVis (Tang et al., 2016) to reduce the embeddings to two dimensions.
+G.2 Indicator Sequence Visualization
+To intuitively examine how FinSeer embeds indicator sequences, we use LargeVis [Tang et al., 2016] to reduce the embeddings to two dimensions.
 In Figure 7, we visualize [^6] https://medviz.org/app/ the vector space of nine indicators, including adjusted close price, open price, high price, low price, close price, volume, movement, returns, and WAP.
@@
 However, this also suggests that deeper relationships between stock movements and indicators remain to be explored, highlighting the potential for further research in this area.
-Figure 3. Visualization of indicator sequence embeddings on ACL18 dataset.
-[…]
-5. Related Work
-5.1. Stock Movement Prediction
-Non-LLM Methods.
-Traditional approaches to stock movement prediction have focused on various aspects of financial data.
-One prominent category of methods analyzes stock price sequences and their corresponding technical indicators (Qin et al., 2017; Feng et al., 2018) to identify patterns in historical data for predicting future movements.
-However, due to the complexity of factors influencing stock prices, subsequent methods have incorporated additional contextual information, such as news articles (Ding et al., 2015; Liu et al., 2018) or social media posts (Xu and Cohen, 2018; Wu et al., 2018).
-Despite efforts, these methods are highly susceptible to noise and struggle to analyze the vast and diverse nature of financial information effectively.
-LLM-based Methods.
-Recent studies have explored using LLMs for financial prediction tasks, either by fine-tuning open-source models or prompting advanced models like GPT-4.
-However, even state-of-the-art models, including GPT-4, have achieved only random-guessing-level accuracy in stock movement prediction (Xie et al., 2023b, 2024a, a).
-This highlights the inherent challenges in identifying and analyzing meaningful patterns in a domain as volatile and multifaceted as stock market prediction.
-5.2. Time-series forecasting with LLMs
-Non-retrieval Methods.
-To enhance the performance of LLMs in time-series forecasting, existing methods have primarily focused on aligning temporal and textual data, either by transforming time-series into textual formats or by encoding both modalities into a unified vector space.
-For instance, a study (Jin et al., 2023) reprograms time-series data into textual representations suitable for LLMs, enhancing prediction accuracy via declarative prompts.
-Similarly, some other studies (Yu et al., 2023; Liu et al., 2024) explore cross-modal alignment: the former applies LLMs to financial forecasting by integrating stock prices with news data, while the latter introduces a cross-modality framework to align time-series with text for improved predictive performance.
-While these advancements improve alignment, existing methods often struggle to process large-scale time-series data comprehensively due to LLM input limitations.
-As a result, effectively leveraging extensive historical data remains a challenge.
-This highlights the need for retrieval-augmented methods, a gap our approach directly addresses.
-RAG Methods.
-With the rapid advancement of RAG techniques in various applications of LLMs, recent research has begun exploring RAG for time-series forecasting.
-For instance, TimeRAG (Yang et al., 2024) integrates RAG into time-series forecasting by combining Dynamic Time Warping (DTW) with LLMs to improve prediction accuracy.
-However, relying solely on numeric similarity is insufficient for financial time-series forecasting, as it fails to capture deeper semantic relationships.
-This underscores the need for a more targeted retrieval framework tailored to the complexities of financial data, a challenge our framework effectively addresses.
-6. Conclusion
-In this work, we introduce the first RAG framework for financial time-series forecasting, featuring a novel retrieval mechanism with three key innovations: a fine-tuned 1B parameter LLM (StockLLM), a candidate selection method leveraging LLM feedback, and a training objective that maximizes similarity between queries and historically significant sequences.
-We also construct new datasets integrating financial indicators and historical stock prices, selected through expert consultation and mutual information analysis, to train the retriever and ensure robust evaluation.
-Our retriever, FinSeer, leverages these datasets and LLM feedback to uncover deeper patterns beyond numeric similarities.
-Experiments demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, while FinSeer surpasses existing retrieval methods, achieving an 8% higher accuracy on BIGDATA22 and retrieving more impactful sequences.
-This work highlights the importance of tailored retrieval models in financial forecasting and provides a robust framework for future research.
-Appendix A Data examples
-A.1. Examples in the Datastore
-An example in our datastore is shown below.
-"stock_name":"ABBV",
-"query_date":"2014-06-05",
-"query_movement":"rise",
-"open":54.549999,
-"high":55.32,
-"low":54.360001,
-"close":55.299999,
-"adj_close":36.980961,
-"volume":4847300,
-"MACD_Histogram":0.0621141602,
-"macd_crossover":null,
-"bollinger_bands":null,
-"exceeding_upper":null,
-"exceeding_lower":null,
-"overbought_and_oversold_conditions":null,
-"kdj_crossover":null,
-"Returns":0.0131913094,
-"VWAP":54.9933333333,
-"alpha_smr":0.000549858,
-"alpha_mom":0.0121661224,
-"alpha002":null,
-"alpha006":null,
-"alpha009":-0.481476,
-"alpha012":-0.481476,
-"alpha021":1,
-"alpha023":0.0,
-"alpha024":-0.769035,
-"alpha028":null,
-"alpha032":null,
-"alpha041":-0.1554335256,
-"alpha046":-0.481476,
-"alpha049":-0.481476,
-"alpha051":-0.481476,
-"alpha053":null,
-"alpha054":126.4264005641,
-"alpha101":-18.282056485
-Appendix B Addtional Experimental Results
-B.1. Score Calculation of Financial Indicators
-We compute the MI scores for each indicator using nonparametric estimation methods provided by mutual_info_regression from the Scikit-learn library.
-Then we normalize them into range $(0,1)$.
-These scores provide insight into the strength of the dependency between each financial indicator and the forward return.
-Higher MI scores indicate stronger relationships, highlighting which indicators are most predictive of future price movements.
-We select the top-20 indicator groups with the highest MI scores in our candidates, as is shown in Figure 4.
-Figure 4. Mutual Information(MI) scores between indicators and forward return.
-B.2. Candidate Movement Correlation
-For each query with retrieved candidates, we convert the movements of retrieved sequences and LLM’s prediction to numerical representations.
-${M_{c_{j}},M_{q}}&=\begin{cases}\text{1},&\text{rise}\\ \text{0},&\text{freeze}\\ \text{-1},&\text{fall}\end{cases}\end{aligned}$
-Then, we calculate the average value of the $M_{c_{i}}$ from the five retrieved sequences to represent the RAG-provided result, and the final prediction of the query is recorded as another numerical value:
-[…]
-Lastly, we adopt the Pearson correlation coefficient to compute the correlation:
-[…]
+Figure 7: Visualization of indicator sequence embeddings on ACL18 dataset.
+G.3 Results of StockLLM-8B-Instruct
+We additionally trained StockLLM-8B-Instruct based on LLaMA-3.1-8B-Instruct to evaluate FinSeer’s performance on larger LLMs.
+As shown in Table 6, the results demonstrate that FinSeer can consistently enhance LLM performance regardless of the backbone model’s inherent capability for stock movement prediction.
+Compared to the 1B version, we observe greater performance variability in the bare LLM, while different retrieval methods maintain similar relative impacts on StockLLM’s performance.
+Notably, random retrieval and other retrieval methods continue to show inconsistent effects, sometimes improving but other times degrading performance..
+FinSeer not only outperforms all retrieval-based approaches but also delivers stable performance enhancements to StockLLM, validating the effectiveness of our RAG framework.
+Table 6: Results of stock movement predictions using StockLLM-8B-Instruct and retrieval models.
+Retrieving Methods ACL18 BIGDATA22 STOCK23
+(+ StockLLM-8B-Instruct) ACC MCC ACC MCC ACC MCC
+w/o Retrieval 0.479 -0.041 0.446 -0.108 0.510 0.020
+Random Retrieval 0.509 0.028 0.450 -0.149 0.495 -0.012
+DTW 0.509 0.020 0.457 -0.087 0.496 -0.008
+Instructor 0.497 -0.005 0.457 -0.087 0.506 0.011
+UAE 0.509 0.019 0.465 -0.081 0.509 0.018
+E5 0.484 -0.031 0.453 -0.099 0.498 -0.005
+BGE 0.522 0.058 0.460 -0.104 0.507 -0.001
+LLM Embedder 0.486 -0.034 0.454 -0.123 0.499 -0.002
+FinSeer 0.554 0.122 0.469 -0.065 0.511 0.023